### v1

# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

Abstract.
The increasing prevalence of surveillance cameras in smart cities, coupled with the surge of online video applications, has heightened concerns regarding public security and privacy protection, which propelled automated Video Anomaly Detection (VAD) into a fundamental research task within the Artificial Intelligence (AI) community. With the advancements in deep learning and edge computing, VAD has made significant progress and advances synergized with emerging applications in smart cities and video internet, which has moved beyond the conventional research scope of algorithm engineering to deployable Networking Systems for VAD (NSVAD), a practical hotspot for intersection exploration in the AI, IoVT, and computing fields. In this article, we delineate the foundational assumptions, learning frameworks, and applicable scenarios of various deep learning-driven VAD routes, offering an exhaustive tutorial for novices in NSVAD. This article elucidates core concepts by reviewing recent advances and typical solutions, and aggregating available research resources (e.g., literatures, code, tools, and workshops) accessible at https://github.com/fdjingliu/NSVAD. Additionally, we showcase our latest NSVAD research in industrial IoT and smart cities, along with an end-cloud collaborative architecture for deployable NSVAD to further elucidate its potential scope of research and application. Lastly, this article projects future development trends and discusses how the integration of AI and computing technologies can address existing research challenges and promote open opportunities, serving as an insightful guide for prospective researchers and engineers.
1. Introduction
As one of the core technologies of the ubiquitous Internet of Video Things (IoVT), Video Anomaly Detection (VAD) aims to use video sensors to automatically discover unexpected spatial-temporal patterns and detect unusual events that may cause security problems or economic losses, such as traffic accidents, violent behaviors, and offending contents (Liu et al., 2024c). With the widespread use of surveillance cameras in smart cities (Ramachandra et al., 2020) and the boom of online video applications powered by 4/5G communication technologies, traditional human inspection is no longer able to accurately monitor the video data generated around the clock, which is not only time-consuming and labor-intensive but also poses the risk of leaking important information (e.g., biometrics and sensitive speech). In contrast, VAD-empowered IoVT applications (Jedari et al., 2020), such as Intelligent Surveillance Systems (IVSS) and automated content analysis platforms, can process massive video streams online and detect events of interest in real-time, sending only noteworthy anomaly parts for human review, significantly reducing data storage and communication costs, and helping to eliminate public concerns about data security and privacy protection. As a result, VAD has gained widespread attention in academia and industry over the last decade and has been used in emerging fields such as information forensics (Yu et al., 2023), industrial manufacturing (Liu et al., 2024a) in smart cities as well as online content analysis in mobile video applications (Yin et al., 2024).
VAD extends the data scope of conventional Anomaly Detection (AD) from time series, images, and graphs to video, which not only needs to cope with the endogenous data complexity, but also needs to take into account the computational and communication costs in resource-limited devices (Jiang et al., 2021). Specifically, the inherent high-dimensional structure of video data, high information density and redundancy, heterogeneity of temporal and spatial patterns, and feature entanglement between foreground targets and background scenes make VAD more challenging than traditional AD tasks at the levels of representation learning and anomaly discrimination (Mishra et al., 2024). Existing studies (Liu et al., 2018b, 2024b; Bao et al., 2022; Li et al., 2024) have shown that high-performance VAD models need to target the modeling of appearance and motion information, i.e., the difference between regular events and anomalous examples in both spatial and temporal dimensions. In contrast to time series AD that mainly measures periodic temporal patterns of variables, and image AD which only focusing on spatial contextual deviations, VAD needs to extract both discriminative spatial and temporal features from a large amount of redundant information (e.g., repetitive temporal contexts and label-independent data distributions), as well as to learn the differences between normal and anomalous events in terms of their local appearances and global motions (Qiu et al., 2024).
However, video anomalies are ambiguous and subjective (Hojjati et al., 2024). The same behavior may be categorized as different types in different scenarios and settings. For example, riding a horse in a grassland is usually normal, whereas a horse appearing on a highway would be considered an anomaly. On the one hand, compared to regular events, anomalies in the real world are difficult to be comprehensively predefined and have a much lower overall frequency of occurrence, making them difficult to collect. Labeling a sufficient number of all possible abnormal samples for model training is almost impossible. As a result, traditional supervised learning-based classification models are usually ineffective in dealing with AD tasks (Mašková et al., 2024). On the other hand, since video storage and transmission costs are significantly higher than other data modalities, engineers favor processing such data on the end-side or edge-side to reduce communication overhead. As we all know, such devices, including surveillance cameras, smartphones, and local servers, are computationally resource-limited. Therefore, it is highly practical to develop deployable VAD systems for real-world applications, which requires concerted efforts by AI beyond communities.
In this article, we extend the conventional scope of VAD from algorithm engineering on spatial-temporal anomaly detection to practical research towards real-world applications, termed Networking Systems for Video Anomaly Detection (NSVAD), to engage a broader readership from the IoT and computing communities. According to research objectives and involved domains, NSVAD is delineated into the hierarchical architecture shown in Fig. 1, encompassing: 1) Hardware Layer, consisting of various video sensors, communication units, and computing servers, etc, responsible for data acquisition, transmission, processing, and result reporting, as well as device networking; 2) System Layer, targeting resource optimization and algorithm deployment platforms for large-scale IoT applications, linking the Hardware Layer and the Algorithm Layer, supporting configurable deployment of VAD tasks on various terminals; 3) Algorithm Layer, focusing on the development of detection algorithms and scene-specific models driven by artificial intelligence, especially deep learning; 4) Application Layer, encompassing IVSS in modern factories, agriculture, and smart cities, as well as various online video applications powered mobile internet. Most existing works belong to Algorihm layer and solely concentrate on VAD model design, overlooking resource costs and challenges in real-world scenarios. For large-scale IoT and mobile video internet, NSVAD with stable performance and reasonable overheads that supports online detection necessitates sensor networking research and support for resource optimization from computing communities. We review the recent advancements and typical methods in the algorithm layer and provide our latest explorations in the system layer to inspire readers to develop NSVAD toward real-world scenarios.
Thanks to the development of edge AI (Walia et al., 2023) and artificial neural networks (Li et al., 2021b), deep learning-driven NSVAD algorithms have made significant progress in recent years and derived from Unsupervised VAD (UVAD), Weakly-supervised VAD (WsVAD), and Fully-unsupervised VAD (FuVAD) routes (Liu et al., 2024c). They have liberated human beings from massive videos and alleviated the public information security concerns. Compared to early manual feature engineering, deep architectures such as convolutional neural networks and attention mechanisms can extract spatial-temporal representations from video sequences end-to-end without applying human a priori, empowering IVSS to process videos in different resolutions and scenarios. Therefore, researchers in this field are currently focusing on the deep structure design and optimization leaning strategies. They have creatively proposed multimodal VAD (Wu et al., 2020, 2022; Tian et al., 2021), Open-Set AVD (OSVAD) (Acsintoae et al., 2022; Zhu et al., 2022), Open-Vocabulary VAD (OVVAD) (Wu et al., 2024b), video anomaly segmentation (Tian et al., 2024), and anomaly retrial (Wu et al., 2024a) tasks as well as integrated detection systems that can be deployed in practical scenes, such as modern manufacturing (Liu et al., 2024b), smart city (Liu et al., 2023c), and automated driving (Wang et al., 2021). In addition to algorithm design, researchers from Networking Systems of AI (i.e., research on the deep convergence of communication and AI) (Song et al., 2022) and IoVT (i.e., subfields of IoT focusing on video sensor design, networking, and data processing) (Chen, 2020) have begun to explore the design deployment-oriented VAD systems to collaboratively deal with multiple challenges (e.g., multi-view cross-scene heterogeneous videos and communication overheads) that come from the dynamic scenarios at the application layer and the limited resources at the hardware layer. These explorations and progress greatly expand the research boundaries and application scenarios of VAD, promoting it as an intelligent system science, i.e., NSVAD.
Although there have been some reviews (Pang et al., 2021; Cook et al., 2019; Santhosh et al., 2020; Li et al., 2023; Samaila et al., 2024) focusing on AD and combing its related work, due to the limited research horizon, the early works usually regard VAD as a fringe research task in the AD community. They focus on time series or images but lack an illuminating analysis of the AD task in video data. Recent survey papers (Nayak et al., 2021; Ramachandra et al., 2020; Chandrakala et al., 2022) continue to focus on unsupervised NSVAD routes in the same vein as the conventional AD task, i.e., using only normal samples to train generative models to learn the prototypical patterns of regular events and to discriminate uncharacterizable test samples as anomalies (Luo et al., 2021). Such reviews, while providing a comprehensive taxonomy of VAD research in the outlier detection perspective, have been informative in the last decade or so when unsupervised methods have dominated the NSVAD algorithm research. However, they ignore the emerging weakly-supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Pang et al., 2020; Zaheer et al., 2022) routes, making them of limited value in guiding further research.
Considering the differences in knowledge bases and orientations of readers, this article provides an in-depth analysis of the basic concepts, related knowledge, and recent advances involved in NSVAD research and summarizes the available research resources. We systematically analyze the assumptions, frameworks, scenarios, advantages, and disadvantages of unsupervised, weakly supervised, and fully unsupervised VAD routes and explain in detail the relevant domain knowledge involved in each route. In addition, we introduce our NSVAD systems designed for dynamic environments in complex scenarios such as industrial IoT and smart cities to guide the NSVAD research in specific applications. Finally, we discuss the research challenges, trends, and possible opportunities to inspire future exploration.
1.1. Attention Analysis
We searched the number of publications and citations with the topic of ”video anomaly detection” in various mainstream academic databases (e.g., IEEE Xplore, ACM Digital Library, SpingerLink, ScienceDirect, and DBLP) to quantitatively present the research hotness, as shown in Fig. 2(a). Early VAD works were limited by hand-crafted features, which cannot handle complicated videos and rely on human a priori, thus having a slow start. Encouraged by the remarkable success of deep learning in video understanding tasks (e.g., action recognition, scene understanding, and multimodal perception), VAD research saw a boom after 2010, with an explosive growth in the number of publications and citations that continues to the present. On the one hand, the spread of surveillance cameras and streaming media platforms has provided sufficient data support for NSVAD research, making it possible to train large-scale deep models with high-performance GPUs. On the other hand, the increasing demand for offending video content detection in various scenarios drives many researchers and engineers from AI and IoT fields to devote themselves to NSVAD research.
Fig. 2(b) presents the AD-related workshops in the top computer science conferences in the past four years. The changes in data types and application scenarios in these workshops show that AD tasks on visual data, especially videos, have dominated the community. For example, the ”DeeperAction” workshop explicitly identified anomalous behavior recognition in surveillance videos as the following research hotspot in behavior analysis. The first ”ASTAD” workshop at WACV’24 centered on anomalous detection of spatial-temporal data and its application to computer vision tasks. In addition, the latest workshops, ”BRAVO” and ”VISION”, explored the application of AD technology in areas such as autonomous driving and modern manufacturing, further demonstrating the high hotness and broad application prospects of VAD. To make it easier for beginners to navigate these workshops, we have categorized all the available resources, please see our public GitHub repository at https://github.com/fdjingliu/NSVAD.
1.2. Related Work
| Year | Ref. | Perspective & Main Focus | Type | Research Routes & Open Task | Content Analysis | |||||||
| Survey | Tutorial | UVAD | WsVAD | FuVAD | OSVAD | OVVAD | Advances Review | Basic Knowledge | Practical Cases | |||
| 2018 | (Kiran et al., 2018) | Un- and Semi-supervised VAD | ✓ | ● | ○ | ◐ | ○ | ○ | ● | ○ | ○ | |
| 2019 | (Cook et al., 2019) | Time Series AD in IoT | ✓ | ☉ | ☉ | ☉ | ☉ | ☉ | ● | ● | ○ | |
| 2020 | (Santhosh et al., 2020) | VAD in Traffic scene | ✓ | ● | ◐ | ○ | ○ | ○ | ● | ○ | ○ | |
| 2022 | (Ramachandra et al., 2020) | Unsupervised VAD in single-scene | ✓ | ● | ◐ | ◐ | ○ | ○ | ● | ● | ○ | |
| 2021 | (Nayak et al., 2021) | Deep learning-besdUnsupervised VAD | ✓ | ● | ○ | ○ | ○ | ○ | ● | ○ | ○ | |
| 2021 | (Rezaee et al., 2021) | Unsupervised VAD in crowd sence | ✓ | ● | ○ | ○ | ○ | ○ | ● | ○ | ○ | |
| 2021 | (Pang et al., 2021) | Deep learing-besd Unsupervised AD | ✓ | ☉ | ☉ | ☉ | ☉ | ☉ | ● | ● | ○ | |
| 2021 | (Blázquez-García et al., 2021) | Time Series AD | ✓ | ☉ | ☉ | ☉ | ☉ | ☉ | ● | ○ | ○ | |
| 2022 | (Chandrakala et al., 2022) | Unsupervised and supervised VAD | ✓ | ● | ◐ | ○ | ○ | ○ | ● | ○ | ○ | |
| 2022 | (Raja et al., 2022) | Unsupervised VAD | ✓ | ● | ● | ○ | ○ | ○ | ● | ○ | ○ | |
| 2023 | - | Un- and Weakly-supervised VAD | ✓ | ● | ● | ○ | ○ | ○ | ● | ◐ | ○ | |
| 2024 | (Liu et al., 2024c) | Generalized VAD | ✓ | ● | ● | ● | ◐ | ○ | ● | ○ | ○ | |
| 2024 | Ours | NSVAD routes and cases in video IoT | ✓ | ✓ | ● | ● | ● | ● | ● | ● | ● | ● |
●: Systematic compendium and presentation to the routes/tasks/contents. ◐: Briefly mentioned. ○: Not presented. ☉: Not applicable.
To the best of our knowledge, this article is the first tutorial-type paper on NSVAD, which systematically presents the basics and recent advances in different research VAD routes and provides practical cases. Most previous papers (Liu et al., 2024c; Ramachandra et al., 2020; Chandrakala et al., 2022; Kiran et al., 2018) focus only on literature reviews, and conference tutorials1 lack of systematic content organization, as shown in Table 1. Compared with existing survey papers, we consider the limited knowledge accumulation of non-specialists, focusing on clarifying the basic concepts and explaining the background models. We present the task definitions, basic assumptions, learning frameworks, and practical cases of unsupervised, weakly-supervised, and fully-unsupervised VAD routes and various emerging tasks (e.g., open-set VAD) for the first time.
In the start-up phase, VAD is often regarded as a fringe task in the generalized AD community, which is only briefly mentioned in overviews of image and time series AD (Cook et al., 2019; Pang et al., 2021), leaving a comprehensive literature survey and systematic progress compilation in the air. Fortunately, researchers in recent years have recognized the continued buzz around VAD and attempted to sort out the current state of research in the field, e.g., (Ramachandra et al., 2020; Chandrakala et al., 2022; Raja et al., 2022). However, limited by the research perspectives, such works mainly focuse on unsupervised methods in the same vein as conventional AD tasks and lacks the organization for emerging research routes like WsVAD and FuVAD. On the one hand, weakly supervised and fully unsupervised routes are detached from the basic learning paradigm of unsupervised out-of-distribution detection for traditional AD research, even if WsVAD belongs to close-set instead of open-set tasks. On the other hand, WsVAD and FuVAD were first proposed in 2018 (Sultani et al., 2018) and 2020 (Wu et al., 2020), respectively, so the above survey can’t cover such new breakthroughs due to timeliness reasons. However, due to the reliable results and excellent performance in real-world videos, WsVAD has been recognized as one of the key routes to advance the application of VAD in intelligent surveillance systems. In addition, fully unsupervised methods are entirely free from restrictions on the composition and labeling of training data, making it possible to learn efficient VAD models using unprocessed large-scale video streams from mobile internet. The emergence of new datasets and models suggests that WsVAD and FuVAD have become non-negligible VAD routes. In view of this, the latest survey (Liu et al., 2024c) summarizes all these routes and proposes a generalized video anomaly event detection framework. The authors combed through more than 200 papers, demonstrating the research hotspots of VAD. However, emerging research tasks, such as OSVAD) (Acsintoae et al., 2022; Zhu et al., 2022) and OVVAD (Wu et al., 2024b), fail to be mentioned. Such tasks greatly enhance the applicability of VAD techniques to online videos and complex scenarios.
Researchers in the AD community have paid attention to the continued progress and application prospects of VAD and introduced it to attendees through conference tutorials. For example, Pang et al. organized a tutorial titled ”Recent Advances in Anomaly Detection” at CVPR2023, focusing on recent work on deep learning-driven unsupervised and weakly supervised VAD. We think this is a good start and highly appreciate the contribution of the organizers to the field. However, a more unified conceptual statement, more comprehensive documentation, and more systematic analysis of challenges and opportunities are necessary to inspire a wider community of readers.
1.3. Contribution Summary
Given that NSVAD research has become an explicit hotspot in AI, computing, and IoT communities, and shows great potential for applications in emerging scenarios such as smart cities and mobile internet, we aim to provide a systematic and inspiring guidance. This article is aimed at researchers and engineers who understand the main concepts and basic knowledge of AI but have no experience in NSVAD. We provide a comprehensive statement of unsupervised, weakly supervised, and fully unsupervised VAD routes, as well as various types of emerging tasks to satisfy readers with different backgrounds and needs. The contributions of this article can be summarized in the following four points:
To the best of our knowledge, this article will be the first tutorial-type paper focusing on Networking Systems for Video Anomaly Detection, which not only provides a well-structured guide for non-specialized readers but also promises to bring together researchers from AI, IoT, and computing societies to promote NSVAD research.
Focusing on AD in IoVT from the NSAI perspective, we comprehensively sort out the UVAD, WsVAD, and FuVAD routes and state their basic assumptions, learning paradigms, applicability scenarios of each scheme.
We open source available resources (e.g., benckmark datasets, code bases, literatures, workshops, and tutorials) and provide our studies on NSAVD in industry and smart cities.
We analyze the developmental vein between various research routes by empirically reviewing the recent advances and discussing the future vision of NSVAD in the context of trends and concerns in NSAI and IoVT.
1.4. Section Navigation
Based on the relationship between the content of the individual sections and the NSVAD architecture shown in Fig. 1, we organize the remainder of this article and provide an intuitive navigation map shown in Fig. 3. Specifically, in Section 2, we state the general basics of VAD, including the task definition, type of anomalies, and application areas. Sections 35 elaborate on the learning paradigms and typical models of UVAD, WsVAD and FuVAD, respectively. We provide detailed explanations of classic methods for readers to understand the core ideas and implementations better. To assist readers to conduct research instantly, we provide a comprehensive introduction to existing datasets and evaluation metrics involved in the current work in Section 6. Section 7 presents our latest exploration of NSVAD, including how to develop high-performance VAD models for dynamic environments in industrial IoT and how to address arbitrary data bias in smart city scenarios. Finally, Section 8 discusses the future vision of NSVAD, providing an in-depth analysis of its existing challenges, development trends, and possible opportunities. This tutorial paper is intended for non-specialists, so we prioritize conceptual clarification and research horizon construction in the main papers. Content that may overlap with existing work (e.g., itemized introductions to reviewed papers, detailed explanations of classical methods, and comprehensive presentations of research case) is relegated to the Supplement.
2. Generalized Foundations of NSVAD
As a cross-cutting research topic, NSVAD has received attention from researchers from the fields of deep learning, video surveillance, mobile internet, and edge computing. In the early stage, VAD follows the problem definition and basic setting of conventional AD tasks, i.e., anomalous samples are regarded as outliers with different distributions (Ramachandra et al., 2020). The corresponding unsupervised methods (Cao et al., 2024; Han et al., 2024b; Ye et al., 2024) attempt to learn prototypical representations of regular events and consider test videos outside the distribution as anomalies. Facing the high-dimensional, complex background and entangled representational properties of video, researchers at this time, mainly with AD backgrounds, focused on introducing efficient video representation learning (e.g., Auto-Encoders (AEs) (Park et al., 2020), Generative Adversarial Networks (GANs) (Sun et al., 2024), Transformers (Feng et al., 2021b), and diffusion model (Basak and Gautam, 2024)) from the CV community.
With the emergence of large-scale video datasets and continuously updated high-performance GPUs, deep learning-driven Video Understanding (VU) technicals have made significant progress, facilitating the flow of VAD research to the cross-cutting topics of VU and AD tasks. Eventually, a series of new VAD routes, such as WsVAD with Multiple Instance Learning and FuVAD based on iterative learning. These methods depart from the open-world assumption under the unsupervised AD setting, i.e., that anomalies are varied and unbounded in real-world situations. Specifically, WsVAD introduces anomaly instances in the training set to facilitate the model’s understanding of the essential differences between negative and positive samples. Although WsVAD methods need to collect and label anomalies, abnormal events of interest in a given scenario are usually enumerable. WsVAD can output more reliable results for specific types of abnormal events because anomalous patterns are seen during the learning phase. FuVAD completely escapes the limitation of UVAD and WsVAD on training data and tries to learn anomaly detectors directly from raw videos, saving the data preparation cost and avoiding the negative effect of mislabeling.
In recent years, the rise of multimodal learning (Xu et al., 2023; Ektefaie et al., 2023) and large multimodal models, such as CLIP (Radford et al., 2021; Wu et al., 2024c) and DALL-E (Reddy et al., 2021) has brought a new windfall for VAD. Inspired by this, researchers have proposed new tasks such as open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022) and open-vocabulary VAD (Wu et al., 2024b), indicating the trend of integration between VAD and generative AI research. We understand the cognitive differences resulting from the research backgrounds of these fields and view the emergence of new routes and tasks as a positive signal to promote VAD to systematic NSVAD research. In addition, NSVAD systems deployed in real-world scenarios must face the challenge of domain bias due to multi-view, cross-scenario videos and consider the limited storage and communication resources of end devices. Recent NSVAD advances (Liu et al., 2024b; Wu et al., 2020) have begun considering both algorithm optimization and model deployment to balance performance and overhead. This section aims to provide a unified concept definition for various WSVAD routes.
2.1. Related Terms
We first introduce AD-related terminologies, including Anomaly Detection (AD), Novelty Detection (ND), Open Set Recognition (OSR), and Outlier Detection (OD). It is worth mentioning that these terms are confusing for most researchers from the computer vision and IoT communities, and AD practitioners often struggle with the lack of unified conceptual explanations (Yang et al., 2021). To this regard, we follow the general community consensus and combine our experience to sort out the above terms, as shown in Fig. 4.
AD aims to detect samples that deviate from the learned normality predefined by the training data (Samariya and Thakkar, 2023). Such deviations are usually classified into two categories: 1) covariate shift at the raw data level, i.e., label-independent distributional differences between semantically identical samples, such as visual effects due to differences in image styles, video scenes, and acquisition equipment parameters, and 2) semantic shift at the label level, i.e., samples belonging to different categories. The former, also known as Sensory Anomaly Detection (SenAD), aims at detecting samples with label-independent bias, and tasks associated with it include domain adaptation and domain generalization. Obviously, UVAD aims to detect anomalous samples with different semantic labels from the training set, termed Semantic AD (SemAD),.
ND is often confused with AD tasks due to the fact that its setting and goals are essentially the same as those of SemAD (Roady et al., 2020). Specifically, ND, also known as novelty class detection, aims to detect test samples that do not belong to any of the known training classes, suggesting that it is primarily concerned with semantic offsetting. It is worth noting that although ND is subdivided into single-class ND and multi-class AD based on the number of categories in the training samples, they are both modeled as binary classification problems, i.e., they only need to identify the unknown categories without focusing on the so-called secondary labels (Yang et al., 2021). In our opinion, the difference between ND and SemAD is that the former has an optimistic attitude towards unknown category data and does not regard them as erroneous as SemAD does. Thus, UVAD can also be referred to as video novelty detection.
OSR aims to train a Multi-Calss Classifier (MCC) to simultaneously accomplish fine-grained categorization of the in-distribution data while detecting the unknown data that only appear in the test phase, i.e., it requires traditional MCC to exclude novel categories automatically (Geng et al., 2020). None of the current VAD routes consider the specific categories of normal events, focusing only on how to discriminate abnormal samples. Although the training set usually contains different normal events, they are all simply labeled as 0. Although OSR and VAD are currently distinct in terms of technical solutions and objectives, complex multi-agent systems (e.g., autonomous cars) often require the model not only to detect arbitrary anomalies that have not been seen before but also to accurately categorize regular events for decision-making. The integration of OSR with VAD has given rise to the Open Vocabulary VAD (OVVAD).
OD aims to detect significantly different samples from other observations, i.e., outliers (Boukerche et al., 2020). Compared to AD, ND, and OSR, where out-of-distribution samples are only visible in the testing phase, OD directly accepts all observations containing both in- and out-of-distribution samples as inputs in the training phase, which is the same as FuVAD. Therefore, compared to UVAD following the SemAD/ND task setting, FuVAD does not restrict the training data to include only regular instances but uses unfiltered videos to learn anomaly classifiers directly. Consequently, FuVAD is superior in dealing with the massive real-time video streams generated by large-scale IoVT systems in real world.
2.2. Definition and Type of Anomaly
NSVAD aims to automatically detect unusual events that may result in personal injury or financial loss. How anomalies are defined and what types are included are critical to conducting NSAVD research for real-world applications. As stated, UVAD and FuVAD follow the basic setups of SemAD and OD tasks, respectively. Thus, they define anomalies as a relative concept, i.e., samples distinguished from common data are all anomalies. In contrast, WsVAD departs from the open-set setting and focuses only on pre-defined specific kinds of anomalies, usually behaviors that humans intuitively perceive as abnormal. In Fig. 5, we show the composition of the training data for various routes.
Anomalies can be categorized into sensory anomalies focusing on domain offsets at the raw data level and semantic anomalies manifesting as semantic label distinctions. VAD aims at detecting anomalous events with positive semantic labels and is uninterested in label-independent offsets caused by the scene, acquisition device, and camera angle. Therefore, the anomalies defined in both UVAD and FuVAD fall into the category of semantic anomalies. Further, according to the information dimensional of the anomalies presented, we refer to (Ramachandra et al., 2020) to categorize them into appearance-only, motion-only, and joint appearance-motion anomalies, which correspond to deviation in spatial, temporal, and spatial-temporal interactions, respectively. Taking the CUHK Avenue dataset (Lu et al., 2013) as an example, the collector placed a red bag on the lawn, regarded as an appearance-only anomaly. Motion-only anomalies are usually similar to regular events in appearance but present incomprehensible motion, such as people wandering in public areas and cars traveling against the traffic flow. In the real world, anomalies usually present complexity and entanglement, not as single-dimensional information deviations but as misalignments in spatial-temporal interactions, termed appearance-motion united anomalies. It is difficult to effectively monitor such anomalies by relying on anomalous object detection or motion analysis alone. Therefore, the key to efficient UVAD is to make the model fully understand the prototypical features of regular events in appearance, motion, and spatial-temporal interactions (Liu et al., 2024b). Multi-proxy task-based UVAD models take this concern into account, aiming to utilize multiple optimization objectives to improve the model’s understanding of the difference between regular events and anomalous instances in different information dimensions.
In contrast, anomalies in the WsVAD dataset are more in line with human perceptions and real-world requirements, which are typically hazardous events that may pose a safety risk in a specific environment. For example, UCF-Crime (Sultani et al., 2018) collects 13 categories of common crimes against public safety, including abuse, arrest, arson, assault, road accidents, burglary, explosions, fights, robberies, shootings, theft, shoplifting, and vandalism. The XD-Vioence dataset (Wu et al., 2020) collects video and audio data of abuse, car accidents, explosions, fights, riots, and shootings. Although WsVAD violates open-world assumptions and fails to detect arbitrary anomalies outside of expectations, it has become a hot topic because of its reliable results on anomalies of human interest. Dataset producers usually predefine the anomalies according to the task requirements. Inspired by existing studies, WsVAD anomalies can be classified into short-term, long-term, and crowd anomalies. Compared with event labels, categorization according to the duration and manifestation of anomalies is more in line with WsVAD model development. Specifically, WsVAD mostly follows multiple instance learning, whereas earlier models assumed that the bags from anomaly videos contained only one anomaly segment, which is not valid for long-term anomalies. In addition, the magnitude of pattern differences in the exemplars formed from videos containing crowd anomalies is usually smaller than that of individual anomalies. The above two issues have been recognized as challenges that limit WsVAD (Liu et al., 2022b). In response, researchers have proposed Top- (Feng et al., 2021a), feature magnitude learning (Tian et al., 2021), and other improvement strategies. Therefore, how to make the model accurately correspond to more than one positive instance in long-term anomalies and perceive the weak differences between anomalies and normals in crowd anomalies is the key to developing high-performance WsVAD.
2.3. Definition of Various NSVAD Routes
UVAD refers to NSVAD schemes that use only easily collected routine events to train models to learn the spatial-temporal pattern boundaries of normal samples and to discriminate anomalies by measuring the distance between test samples and learned normality (Ramachandra et al., 2020). UVAD dominated early VAD research because it follows the open-world assumptions in the same vein as the AD community, circumventing predefinitions and collecting anomalous instances (Liu et al., 2021a). In Fig. 6, we show the general learning framework of UVAD and compare it with WsVAD and FuVAD. Specifically, UVAD assumes that models trained on regular events will only describe normal spatial-temporal patterns and will exhibit significant deviations when confronted with unseen anomalous examples, such as probability distributions (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), distances (Dalal et al., 2006; Saligrama and Chen, 2012; Dalal and Triggs, 2005), and proxy task errors (Liu et al., 2022c; Hasan et al., 2016; Gong et al., 2019; Liu et al., 2024b). Early approaches first used local binary operators (Hu et al., 2018; Nawarathna et al., 2014; Zhang et al., 2015), spatial-temporal points of interest (Dollár et al., 2005), etc. to characterize spatial-temporal features that are normal events, and then employed One-Class (OC) classifiers (e.g., OC support vector machines and OC neural networks) (Ionescu et al., 2019; Sun et al., 2019) to learn the pattern boundaries, and considered test samples whose features fell outside the boundaries as anomalous. Such methods rely on manual features and are prone to dimensionality disasters.
In recent years, deep learning-driven UVAD integrates feature extraction and normality learning into a unified framework with two phases, training and testing, which corresponds to normality learning via the use of negative samples and anomaly detection via the detection of out-of-distribution samples (Liu et al., 2024c). In the normality learning phase, the network learnable parameters are optimized by minimizing a loss function over all negative samples. Whereas, in the testing phase, the degree of anomaly is measured by quantifying the distance between the test samples and the learned normality. Among them, reconstruction-based methods have dominated UVAD research in recent years (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020; Liu et al., 2022c, 2024b). On the one hand, most of the challenges faced by such methods, such as global motion modeling, temporal normality learning, and spatial detail inference, have been intensively studied in video self-supervised learning. As a result, many methods have driven the development of UVAD by drawing inspiration from existing methods such as video prediction (Oprea et al., 2020; Wang et al., 2022). On the other hand, since reconstruction/prediction methods aim to learn a generative model that can reason about regular events, their basic settings and optimization goals are clear and unambiguous, and thus easy to implement and follow. Essentially, UVAD is transductive learning, i.e., model training and testing are relatively independent. Numerous studies have shown that the performance of models in the normality learning phase on the agent task does not show a positive correlation with downstream anomaly detection. Due to the diversity of events, the spatial-temporal features of normal and abnormal samples overlap, and UVAD usually fails to actively recognize discriminative features and incorrectly learns some shared patterns when only regular events are available for model training (Liu et al., 2023c). In addition, Park et al. (Park et al., 2020) pointed out that overpowered deep neural networks may be able to effectively reason about unseen anomalous events during the testing phase due to overgeneralization performance, which may lead to underdetection. In response, the researchers proposed a memory network enhancement approach to weaken the model’s ability to generalize representations of anomalies by recording prototypical features.
WsVAD uses weakly semantic video-level labels to supervise the output of strong semantic frame-level labels, i.e., frame-by-frame anomaly scores, by the sequencing model, thus enabling temporal localization of anomalous events. The first weakly-supervised approach is the multiple-example ranking framework proposed by Sultani et al. (Sultani et al., 2018) in 2018, which lays out the basic MIL architecture of the WsVAD route, and whose concurrently publicly available UCF-Crime dataset has become the most widely used weakly-supervised benchmark. They consider a video as a collection of multiple examples (video clips), where clips containing abnormal frames are positive examples, while clips with all normal frames are labeled as negative examples. Obviously, a normal video with label 0 produces all examples called negative bags, and the example-level labels are all . An abnormal video with the label constitutes a positive packet, which contains both positive and negative instances. Inspired by multiple-example learning, WsVAD aims to train a scoring model to output the anomaly scores of each example using video-level labels. The authors introduce a MIL ranking loss inspired by the hinge loss, which encourages the model to output high anomaly scores close to 1 for anomalous clips by maximizing the difference between the anomaly scores of the largest-scoring instances in the positive and negative bags while scoring regular clips as close to 0 as possible. In fact, WsVAD does not belong to any of the classes of out-of-distribution detection tasks introduced in Section 2.1, but is rather a type of multiple instance learning under weak semantic labeling supervision.
Compared to UVAD, the weakly supervised approach introduces anomalous videos in the training set and provides video-level labels for all training samples. Although anomalous examples are diverse and unenumerable in the real world, noteworthy anomalous events in specific scenarios are usually limited and easy to collect, such as thefts, robberies, and traffic accidents, which can be obtained in large quantities from surveillance IoT systems and online video platforms. In addition, the labor cost of video-level tagging, which only requires marking whether a video contains an anomaly without worrying about the specific timing location (frame-level labeling) and detailed anomaly category, is usually affordable. For example, the UCF-Crime dataset is much larger than the one used by UVAD, but only 2100 discrete labels need to be provided. Due to the introduction of an additional human prior and the fact that the model has seen anomalous events during the training phase, WsVAD results are typically more reliable than UVAD, achieving excellent and consistent performance in detecting specific anomalous events. As a result, WsVAD has become a mainstream VAD scheme and is considered as the most promising research route for deployment in intelligent surveillance systems. The latest research attempts to mine anomaly-related clues from audio or subtitle text accompanying video frames, and proposes the multimodal WsVAD (Yang et al., 2024). This article introduces the recent progress of WsVAD in Section 4.
FuVAD attempts to learn anomaly classifiers directly from large-scale raw videos without any editing and labeling (Pang et al., 2020; Zaheer et al., 2022). Specifically, FuVAD’s training data contains both positive and negative samples, and due to the low frequency of anomalous examples compared to regular events, the anomalous samples to be detected can be regarded as outliers with different patterns from the main data. In essence, the FuVAD model is transductive learning and does not follow the training-testing process of UVAD. Fully unsupervised methods have become a research hotspot in the internet era, where data preparation is costly, by virtue of the fact that they do not require any constraints and assumptions on the training data, and can be used to train the model by directly accessing huge amounts of videos from the real world (Liu et al., 2024c).
2.4. Emerging Research Tasks
In Section 2.1, we mentioned one of the inherent drawbacks of WsVAD, which violates the open-set problem property of the AD task by only detecting specific anomalies predefined in the training set and not being able to cope with diverse and arbitrary anomalous events in the open world. Action Recognition (AR) (Kong and Fu, 2022; Sun et al., 2022) and VAD aim to understand specific behaviors. They can learn from each other’s research ideas in data modality and feature learning, and the pre-trained AR models on large-scale video datasets are expected to be directly used for feature extraction and anomaly semantic cue mining in VAD. However, AR follows the closed-set task setting, which can only model the spatial-temporal patterns of known categories of behaviors but cannot empower the model to recognize out-of-distribution samples. In response, researchers have proposed various open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022) schemes to break through the above barriers, as shown in Fig. 7(b).
For example, OpenVAD in (Zhu et al., 2022) aims to integrate the advantages of UVAD, which can handle arbitrary anomalous events, and WAED, which has a low false alarm rate in detecting specific anomalies. The proposed method integrates deep evidence learning and normalized flow into the MIL to equip WsVAD with the ability to identify unknown anomalies by quantifying uncertainty. Acsintoae et al. (Acsintoae et al., 2022) propose a dataset for supervised OSVAD, named UBnormal, that maintains the task’s open-set properties. Since the anomalous behavior in this dataset is generated through the VAD engine, it comes with fine pixel-level labels, making it possible to train VAD models in supervised learning. In short, this dataset attempts to bridge closed supervised learning and open anomaly detection, and experiments show that it can improve performance without compromising the open-set properties of existing VAD models. OVVAD is closest in setting to the OSR task and aims to learn a multi-classifier capable of detecting and classifying all known and unknown anomalous events. Compared to OSVAD, OVVAD is more in line with the display requirements of scenarios such as autonomous driving. The first OVVAD model proposed by Wu et al. (Wu et al., 2024b) splits the task into two complementary tasks, i.e., AD and anomaly classification, and jointly optimizes them using the knowledge from the large models.
2.5. Applications
IoVT (Chen, 2020) has been widely used in daily life so that VAD models can be directly deployed in off-the-shelf cameras without additional hardware costs to realize contactless real-time environmental monitoring of factories, roads, households, and public places. In addition, 5G communication technology has driven the rapid development of mobile internet, and the number of short videos uploaded to online video-sharing platforms from smartphone terminals alone exceeds 400 million per day, making automation of content regulation imminent. We expect that this article will not only inspire VAD researchers to conduct academic studies but also attract the attention of practitioners in the transportation management, modern industrial, retail commerce, and public service industries.
The application scenarios of NSVAD include: a) Traffic accident detection and road security monitoring in smart transportation. Such problems have been the focus of WsVAD. The first weakly supervised dataset, UCF-Crime (Sultani et al., 2018), contains such traffic anomalies. The subsequently proposed ROADMAP dataset (Wang et al., 2021) is all captured from real traffic scenarios, which creates a benchmark for video-driven spatial-temporal anomaly detection research in open space. b) Accident discovery and production line management in model industry, etc. Video can collect large-scale spatial-temporal information without contact, and has the advantages of high information density, wide perception field of view, stable data quality, and adaptability to extreme environments compared to sensors such as temperature and humidity. Therefore, the use of video cameras to monitor the state of industrial production has gained the attention of researchers. As an example, we propose a multi-branch normality learning framework in (Liu et al., 2024b) to cope with high-motion dynamics and complex appearance patterns in industrial scenes. Experiments conducted on realistic datasets validate the effectiveness of the proposed AMP-Net, which reflects the potential of VAD in industrial scenarios. c) Commercial entities directly facing people, such as retail, often face the risk of economic losses due to theft, robbery, and fraudulent behaviors. Both the UCF-crime (Sultani et al., 2018) and XD-violence (Wu et al., 2020) datasets capture such scenarios of violations, providing benchmarks for building crime detection systems for public commercial areas. In addition, VAD can provide new ideas for customer behavior analysis, i.e., detecting abnormal customer flows to guide the adjustment of store layout and business strategies, etc. d) The public service usually needs to bear data privacy concerns, especially for surveillance videos containing biological information (e.g., faces, body behaviors, and sounds). Human inspection is not only time-consuming and laborious but also may cause sensitive information leakage. In contrast, VAD can analyze video on the user side and transfer only the segments of concern to the third parties.
3. Unsupervised Video Anomaly Detection
UVAD follows the general setup of semantic anomaly detection tasks, where only easily collectible regular instances are used to train models describing the normality of videos, detecting anomalies by measuring deviations between test samples and the learned model. From traditional machine learning to deep representation learning, UVAD has undergone multiple revolutions in feature extraction and normality learning, leading VAD to become a key issue in the AD and CV communities. With its close ties to the AD community and its long-standing development history, UVAD has long been regarded by researchers as the mainstream research route of NSVAD algorithms. As a result, existing surveys (Nayak et al., 2021; Ramachandra et al., 2020) typically focus on reviewing UVAD literature, lacking in-depth discussion of emerging weakly supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Zaheer et al., 2022; Pang et al., 2020) routes, and overlooking novel tasks such as open-set and open-word detection. This article not only systematically reviews the latest developments in these new routes and tasks but also provides foundational knowledge and timely and classic methods of UVAD in this section.
Based on the means of normality learning and deviation calculation principles, existing deep learning methods are generally divided into three categories: distance-based (Saligrama and Chen, 2012; Dalal et al., 2006; Dalal and Triggs, 2005), probability-based (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), and reconstruction-based (Liu et al., 2024b, 2018b; Bao et al., 2022; Luo et al., 2017). From our perspective, distance-based UVAD methods are a more general form of probability-based and reconstruction-based methods because probability deviation and reconstruction error calculation are fundamentally just different distance measurement methods. Specifically, distance-based methods include using single-classifier learning of video sequence spatial-temporal representations in the deep feature space, such as OC-SVM and OC-NN (Ionescu et al., 2019; Sun et al., 2019). The drawback of such methods is that the trained models cannot be incrementally compatible with new data, leading to the need to retrain classifiers from scratch when new data is generated, such as in scenarios involving scene transitions. Another type of distance-based method is to use Gaussian mixture models (Wu et al., 2010) to model video feature normal vectors and measure deviations using Mahalanobis distance. In contrast, probability-based methods attempt to map the spatial-temporal representations of regular events into a probability framework and discriminate anomaly instances by measuring differences in probability distributions. Such methods tend to use traditional models such as Markov random fields (Benezeth et al., 2009) to build probability space. Deep learning-based attempts have encountered significant increases in computational costs and are noticeably slower in inference speeds.
In fact, the most prevailing deep UVAD approach is reconstruction-based (Ramachandra et al., 2020). On the one hand, reconstruction-based methods (Liu et al., 2018b; Park et al., 2020; Liu et al., 2024b) aim to train models to represent the general spatial-temporal patterns of regular events through self-supervised proxy tasks, benefiting from advances in video self-supervised learning (Jing and Tian, 2020) and deep neural networks. On the other hand, such methods avoid complex mathematical computations, which are easy to implement and exhibit excellent performance, thus being widely praised by existing researchers. The premise of reconstruction-based methods is that generative models trained on massive normal samples can effectively infer the spatial-temporal patterns of regular events. For anomalous events, the performance of proxy tasks will significantly decrease, and the resulting error can be used as a quantitative basis for measuring deviations to calculate anomaly scores. Common generative models include deep autoencoders (Hasan et al., 2016; Chong and Tay, 2017; Liu et al., 2022a, c; Park et al., 2020), variational autoencoders (Fan et al., 2020; Yan et al., 2018), and generative adversarial networks (Chen et al., 2021; Huang et al., 2021; Nguyen et al., 2020; Li et al., 2022; Cai et al., 2021a; Zhang et al., 2022). Proxy tasks include reconstructing input sequences and predicting future frames, which belong to pixel-level image generation. However, most reconstruction-based methods only calculate frame-level errors as anomaly scores without performing spatial localization, as the spatial contribution of anomalies is typically not significant. Spatial localization can only serve as a quantitative visualization results.
3.1. Taxonomy and Advances
This article also systematically reviews deep learning-driven UVAD methods, providing a taxonomy that can aid in understanding the current state of research and inspire further exploration. However, due to space constraints and the emphasis of the tutorial paper on guiding beginners, we only present the underlying logic of the proposed taxonomy and summarize the research trends of the latest advances in the main text message. For more detailed description of the existing UVAD methods, please refer to Section 1.1 of Supplement.
The methods based on distance and probability described earlier focus more on handcrafted features (Cheng et al., 2015; Ravanbakhsh et al., 2018; Shi et al., 2021; Ravanbakhsh et al., 2017) and traditional classification (Tran and Hogg, 2017; Turchini et al., 2017; Vu et al., 2017) models, which have been primarily surpassed by distance-based methods in the era of deep learning. Therefore, existing classification systems appear outdated in delineating recent advancements and reflecting research trends, failing to highlight the latest challenges and directions. To address this, inspired by data preprocessing techniques and forms of deep learning modeling, we categorize UVAD into two main classes: Global Normality Learning (GNL) and Local Prototype Modeling (LPM), as presented in Table. 2.
GNL utilizes the entire video sequence as input, often requiring no additional preprocessing such as spatial-temporal cube partitioning or foreground object extraction, and employs end-to-end deep neural networks to directly learn video ecologies (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020; Chang et al., 2020; Liu et al., 2024b, 2023c). Over the years, researchers have believed that videos possess two informational dimensions, namely spatial and temporal, corresponding to appearance and motion, requiring different approaches to capture their normality. Therefore, within our UVAD taxonomy, GNL is further subdivided into Single-Proxy Task (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020) and Multi-Proxy Task methods (Chang et al., 2020; Liu et al., 2022c, 2024b). In contrast, LPM methods argue that video data contain a plethora of redundant information tantamount to clues related to normality. Thus, they opt to use spatial-temporal cubes (Mahadevan et al., 2010; Roshtkhari and Levine, 2013; Sabokrou et al., 2017; Liu et al., 2021a) or foreground objects (Bao et al., 2022; Liu et al., 2023a) containing dense and effective information as network inputs instead of the entire video sequence, focusing on prototype feature learning of local patch. We classify such methods into spatial-temporal Patch-based and Object-driven methods.
3.1.1. Global Normality Learning
Convolutional neural network-driven deep representation models can directly learn task-relevant spatial-temporal representations from raw video sequences and can adapt to different scale inputs and feature dimension requirements through simple structural adjustments. Global normality learning aims to learn video normality directly from complete RGB videos or optical flow sequences. Compared to local prototype modeling, GNL requires no additional data preparation and is easy to optimize.
The earliest methods did not distinguish between spatial and temporal information, typically using only RGB videos as input and employing a single self-supervised proxy task (e.g., sequence reconstruction or future frame prediction) to broadly learn spatial-temporal normality. Generally, methods based on a single proxy task focus on designing more efficient single-stream end-to-end deep structures. Recent efforts include the introduction of 3D convolutional networks (Gong et al., 2019) and convolutional long short-term memory networks (Luo et al., 2017) to enhance the representation capability of spatial-temporal features. Subsequent researchers discovered that spatial and temporal information have different characteristics, with the former focusing on local pixel inference and the latter on modeling global dynamics. Moreover, the addition of proxy tasks such as reconstruction and prediction losses often brings additional performance gains without significantly increasing training costs. Thus, they proposed introducing additional proxy tasks within the GNL framework. In addition to the separation of spatial-temporal normality learning (Liu et al., 2022c; Chang et al., 2021) and simple proxy task stacking (Zhao et al., 2017; Chen et al., 2020), recent work has also explored novel tasks such as appearance-motion consistency (Cai et al., 2021b), spatial-temporal coherence (Cheng et al., 2023; Ning et al., 2024), and correlation (Zhao et al., 2022).
3.1.2. Local Prototype Modeling
In contrast, local prototype modeling (Mahadevan et al., 2010; Roshtkhari and Levine, 2013; Sabokrou et al., 2017; Liu et al., 2021a; Bao et al., 2022; Liu et al., 2023a) treats a video as an information cube with dimensions , where and denote the spatial height and width, and represents the number of frames. It is observed that background information repeats across frames. On one hand, anomalies of interest typically occupy only a small portion of the information volume within the entire cube, and direct learning from the complete sequence often entails high computational costs. On the other hand, separating regions with different information densities and modeling their relationships with each other is beneficial for understanding the interaction between events. To address this, researchers have proposed the method of local prototype modeling, aiming to mitigate the handling of repetitive information to reduce training costs and model the relationship between foreground targets and background scenes to enhance anomaly detection performance. According to the data preprocessing methods, we categorize such methods into those based on spatial-temporal Patch-based (STP) and Foreground Object-Driven (FOD). The former employs simple spatial-temporal segmentation to divide the video into several information bodies, while the latter relies on pre-trained object detection models (e.g., RCNN (Girshick, 2015), FPN (Xu et al., 2019), and YOLO (Redmon et al., 2016)) to selectively learn the spatial-temporal normality of specific subjects.
3.2. Classic UVAD Models
Considering the limited guiding value of a mere progress review for beginners, this article has chosen representative methods of UVAD to provide a detailed introduction to the research motivations and core ideas. The selected methods include: 1) Future Frame Prediction (FFP) framework (Liu et al., 2018b), which introduces the video prediction into the VAD task for the first time. 2) Memory-Guided Normality Learning (MGNL) (Park et al., 2020), the first memory network for VAD. We elaborate on the implementation process and comprehensively review the related basic knowledge in Section 2.1 of Supplement.
4. Weakly-supervised Video Anomaly Detection
Inspired by multiple instance learning (MIL) (Herrera et al., 2016), WsVAD organizes videos into bags containing several instances. All segment instances from the same bag share a video-level label. Here, indicates negative bags while indicates positive ones containing at least one anomalous instance. WsVAD strikes a balance between performance and data preparation cost, bridging the gap between UVAD with unreliable results and supervised learning with fine labels. Specifically, unlike unsupervised methods that train models only on regular events and simplely consider all unseen samples as anomalies, WsVAD is trained with positive samples, enabling it to effectively understand the inherent differences between normal and anomalous instances. Consequently, existing research indicates that WsVAD yields more reliable results and outperforms UVAD schemes with lower false alarm rates in real-world tasks, such as crime behavior identification and traffic accident detection. Annotating long video sequences frame by frame for supervised learning is often impractical. For instance, the training set of UCF-Crime dataset includes 1,610 long videos containing 13,741,393 frames in total. In contrast, WsVAD only requires coarse semantic video-level annotations, with each video needing only a discrete binary label. Various types of anomalies are roughly labeled as 1, requiring minimal expert involvement in data annotation and significantly reducing labor costs compared to fine-grained labeling. Thus, although WsVAD no longer adheres to the open-world task setting of the AD community and can only detect specific anomalies, it has garnered widespread attention in recent years due to its stable performance and reliable results. Subsequent works typically adopt the MIL regression task setting proposed by Sultani et al. (Sultani et al., 2018) and utilize their concurrently open-sourced UCF-Crime dataset as a benchmark. To further validate the generalization ability of weakly-supervised models, some researchers relocate positive samples from the test set of UVAD datasets to the training sets and provide video-level labels, proposing reconfigured datasets such as UCSD Ped2 and ShanghaiTech Weakly for WsVAD validation.
4.1. Taxonomy and Advances
The superior performance of WsVAD on real-world videos has inspired researchers to develop VAD models tailored to the complexities of real-life scenarios. In 2020, Wu et al. (Wu et al., 2020) extended VAD research from single-modal video pattern analysis to multi-modal learning to leverage heterogeneous data in real-world scenes for enhanced anomaly event detection. Their collected VD-violence dataset is the first multi-modal VAD benchmark, comprising RGB images and audio modalities, focusing on violence behavior detection in complex scenes. This work not only provides the first multi-modal VAD dataset and solution but also motivates researchers in the community to explore anomaly clues from multi-modal data such as audio and text, leading to a new wave of VAD development: multi-modal WsVAD. In this section, we comprehensively review the latest advancements in both single-modal and multi-modal WsVAD, and illustrate the foundational knowledge and specific implementations required for WsVAD using the Multiple Instance Regression (MIR) framework and the Local-Global Network (LGN) as examples.
4.1.1. Unimodal Methods
Unimodal WsVAD (UWsVAD) methods focus on extracting anomaly-related cues from RGB image sequences. These methods generally follow a three-step process: 1) preprocessing videos into several non-overlapping segments, 2) extracting spatial-temporal features using models like Convolutional 3D Networks (Tran et al., 2015) and Inflated 3D Networks (Carreira and Zisserman, 2017), and 3) computing anomaly scores using a multi-instance ranking loss to differentiate between normal and anomalous instances. The Multi-Instance Ranking (MIR) framework (Sultani et al., 2018) first introduced MIL to WsVAD with a focus on predicting higher anomaly scores for anomalous segments while minimizing score fluctuations. For more detailed description of the existing UWsVAD methods, please refer to Section 1.2.1 of Supplement.
4.1.2. Multi-modal Methods
Multi-modal approaches in WsVAD combine video with audio and other data types. These methods face challenges due to limited benchmark datasets and comparison standards, which has hindered widespread adoption. A key milestone in this domain is HL-Net (Wu et al., 2020), introduced in 2020, marking a significant contribution to multi-modal VAD research with the open-sourced XD-Violence dataset. This dataset has spurred further research by providing a benckmarks for comparisons. Existing multi-modal research mainly explores the fusion of video, audio, and optical flow, as stated in Section 1.2.2 of Supplement.
4.2. Classic WsVAD Models
We select two representative methods from unimodal and multimodal WsVAD, i.e., Multi-Instance Ranking (MIR) (Sultani et al., 2018) and Holistic and Localized Network (HL-Net) (Wu et al., 2020), to elaborate on the concepts of MIL and multimodal information processing in WsVAD research, respectively. Specifically, MIR has laid the foundation for the MIL-based solution, marking a milestone in VAD research. However, its optimization objectives and detailed implementation are challenging for many researchers in the AD and CV communities to understand. Therefore, we provide a comprehensive exposition to the motivation, theoretical logic, and related knowledge. In contrast, HL-Net is the first model for multimodal violence detection. The simultaneous released XD-violence dataset has inspired lots of researchers from the AD and multimodal understanding fields to delve into the emerging hotspot of multimodal VAD. We choose this method as a case study to disseminate knowledge of multimodal understanding to the AD community, aiming to propel multimodal VAD from simple-modal fusion towards systematic multimodal anomaly clues exploration, please see Section 2.2 of Supplement for more information.
5. Fully unsupervised video anomaly detection
FuVAD follows the transductive learning setup of traditional outlier detection tasks, aiming to directly learn an anomaly classifier from all unfiltered observations to detect samples significantly different from the primary data. Existing methods, partly inspired by time series outlier detection research, utilize deep clustering to identify pattern centers of the data and consider samples far from the learned centers as anomalies. However, the pattern dimension of video data is much higher than that of time series, and due to reasons such as similar environmental backgrounds, normal and abnormal event spatial-temporal patterns often overlap, making it infeasible to determine pattern boundaries through clustering when dealing with complex datasets. In recent years, researchers have proposed FuVAD schemes based on iterative learning, gradually amplifying the pattern differences between anomalous samples and dominant normal data through the cooperation of feature extraction modules and anomaly models. Compared to other VAD research routes, FuVAD does not require filtering and labeling training data, directly utilizing unclipped unlabeled monitoring videos to train models, which aligns with the data state-agnostic condition in online learning.
5.1. Recent Advances
Inspired by unmasking, Liu et al. (Liu et al., 2018a) connect heuristic unmasking with multiple classifier two-sample tests, introducing a history sampling method to enhance testing capabilities in video anomaly detection and a motion feature calculation method for better representation and generalization. Li et al. (Li et al., 2021c) use a distribution clustering to identify anomaly example groups, then train an autoencoder with normal data subsets to learn representations of normalcy, iterating this process to refine the encoder’s ability to describe regular events.
Drawing from the Masked Autoencoder (MAE) (He et al., 2022), the Temporal Masked Auto-Encoder (TMAE) (Hu et al., 2022) aims to learn high-quality representations for anomaly detection by employing a visual transformer for completion tasks on spatial-temporal cubes, recognizing the significance of the temporal dimension in video anomalies. This approach is designed to efficiently complete regular events, highlighting anomalies due to their significant loss during completion.
An end-to-end self-training deep ordinal regression (SDOR) framework (Pang et al., 2020) iteratively learns pseudo-normal and anomaly scores from raw sequences, starting with identifying potential anomaly frames using existing algorithms and employing ResNet50 and neural networks for score computation, leveraging self-training for simultaneous optimization of feature learning and anomaly scoring. Generative Cooperative Learning (GCL) (Zaheer et al., 2022) learns anomaly detectors from unlabeled mixed data by exploiting anomaly events’ low frequency, featuring a generator and discriminator working cooperatively. The focuses on regular event representations and uses negative learning for anomalies, generating pseudo-labels for , which estimates anomaly probabilities to further refine .
5.2. Classic FuVAD Models
We introduce two typical deep learning-based FuVAD schemes in Section 2.3 of Supplement, including 1) Self-trained Deep Ordinal Regression (SDOR) (Pang et al., 2020), which utilizes self-training to jointly optimize feature learning and anomaly scorer, and 2) Generative Cooperative Learning (Zaheer et al., 2022), which leverages the low-frequency nature of real-world anomalies to construct pseudo-labels for FuVAD. SDOR (Pang et al., 2020) is the first deep NSVAD method designed for unfiltered and unlabeled videos, which explicitly points out the limited applicability of UVAD and WsVAD in realist scenarios due to the constraints on data and the performance reliance on feature learning. GCL (Zaheer et al., 2022) introduces negative learning to increase the contrast between regular sequences and potential anomalies, directly learning the differences between anomalies and the majority of samples (normal) through the interplay of generator and discriminator.
6. Model Evaluation
| Labeling | Year | Dataset | #Videos | #Frames | #Scenes | #Classes | #Anomalies | ||||||
| Total | Training | Testing | Total | Training | Testing | Normal | Abnormal | ||||||
| Unsupervised | 2008 | Subway Entrance | - | - | - | 144,250 | 76,543 | 67,797 | 132,138 | 12,112 | 1 | 5 | 51 |
| 2008 | Subway Exit | - | - | - | 64,901 | 22,500 | 42,401 | 60,410 | 4,491 | 1 | 3 | 14 | |
| 2011 | UMN† | - | - | - | 7,741 | - | - | 6,165 | 1,576 | 3 | 3 | 11 | |
| 2013 | UCSD Ped1 | 70 | 34 | 36 | 14,000 | 6,800 | 7,200 | 9,995 | 4,005 | 1 | 5 | 61 | |
| 2013 | UCSD Ped2 | 28 | 16 | 12 | 4,560 | 2,550 | 2,010 | 2,924 | 1,636 | 1 | 5 | 21 | |
| 2013 | CUHK Avenue | 37 | 16 | 21 | 30,652 | 15,328 | 15,324 | 26,832 | 3,820 | 1 | 5 | 77 | |
| 2018 | ShanghaiTech | - | - | - | 317,398 | 274,515 | 42,883 | 300,308 | 17,090 | 13 | 11 | 158 | |
| 2020 | Street Scene | 81 | 46 | 35 | 203,257 | 56,847 | 146,410 | 159,341 | 43,916 | 205 | 17 | 17 | |
| 2023 | NWPU Campus | 547 | 305 | 242 | 1,466,073 | 1,082,014 | 384,059 | 1,400,807 | 65,266 | 43 | 28 | - | |
| Weakly Supervised | 2018 | UCF-Crime | 1,900 | 1,610 | 290 | 13,741,393 | 12,631,211 | 1,110,182 | - | - | - | 13 | 950 |
| 2019 | ShanghaiTech Weakly | 437 | 330 | 107 | - | - | - | - | - | - | 11 | - | |
| 2020 | XD-Violance | 4,754 | - | - | - | - | - | - | - | - | 6 | - | |
| Supervised | 2022 | Ubnormal | 543 | 268 | 211 | 236,902 | 116,087 | 92,640 | 147,887 | 89,015 | 29 | - | 660 |
†The frame rate is set to 15 fps. The Ubnormal contains a validation set with 64 videos totaling 14,237 normal and 13,938 abnormal frames.
In this section, we elaborate on the characteristics of prevailing datasets and common evaluation metrics. Based on the annotations, we categorize existing datasets into unsupervised, weakly supervised, and supervised, as presented in Table 3. Some datasets for emerging WsVAD and FuVAD methods are reconstructed from UVAD, so we provide explanations in corresponding sections, refraining from separate introductions. Additionally, we provide the first systematic WSVAD evaluation system, categorizing existing metrics into accuracy-oriented and cost-oriented.
6.1. Public Datasets
6.1.1. Unsupervised Datasets
Subway (Adam et al., 2008) focuses on abnormal event detection in subway stations, comprising two subsets: Entrance and Exit, with abnormal behaviors depicted as shown in Fig. 8. The entrance is used to detect fare evasion behaviors like jumping over turnstiles. In contrast, abnormal events in the Exit include 1) appearance anomalies, such as cleaners wearing attire different from regular passengers, and 2) motion anomalies, such as people walking in the opposite direction of the exit. This dataset was collected 20 years ago, so its resolution is low. Moreover, due to overly coarse annotations and lack of clear rules defining abnormal events, recent VAD works have discontinued using this dataset. Therefore, we recommend researchers to use larger and more clearly annotated datasets, such as UCSD Ped2, CUHK Avenue, and ShanghaiTech, for validation.
UMN (Cui et al., 2011) consists of 11 long video sequences captured from 3 different scenes. Unlike Subway, this dataset includes behaviors pre-set by the collectors and volunteers acting them out, as shown in Fig. 8. Abnormal events are defined as sudden evacuations caused by panic in crowds, simulating mass anomalies in the real world due to terror attacks and fires. Due to the lack of convincing application scenarios, neglect of diversity, and low frequency of real-world anomalies, this dataset is no longer used in UVAD.
UCSD Pedestrian (Li et al., 2013) is one of the most widely used UVAD benchmarks collected by researchers from the campus of University of California, San Diego. Due to its realistic scenes and clear definition of abnormal events, it is widely used to evaluate models’ potential applications in public safety. According to camera angles, UCSD Pedestrian is divided into Ped1 and Ped2 subsets. The latter is more favored by existing methods because the camera view is parallel to the road direction, eliminating the problem of changing object sizes with spatial changes. Recent methods have achieved remarkable frame-level AUCs exceeding 98% on UCSD Ped2, while the best performance on Ped1 is still below 94%, indicating that performance on Ped2 has saturated. Considering that cross-view videos are inevitable in the real world, we recommend that future researchers consider using Ped1 for validating and empowering UVAD models to overcome perspective issues. Some researchers move abnormal samples from the test set of UCSD Ped2 to the training set to meet the requirements of weakly supervised training, proposing a reconstructed UCSD Ped2 dataset for WsVAD testing.
CUHK Avenue (Lu et al., 2013) is similar to UCSD Pedestrian in scene selection and data structures, with only the test set containing 47 abnormal events captured from a university campus. However, CUHK Avenue considers more realistic situations. Collectors collected appearance-only, motion-only, and appearance-motion united anomalies to comprehensively simulate the diversity and complexity of real-world anomalies. Therefore, the appearance and motion patterns of videos in CUHK Avenue are more complex, making it one of the most challenging UVAD benchmarks. Additionally, CUHK Avenue inspired later researchers to collect UVAD datasets by simulating abnormal behaviors, such as ShanghaiTech (Liu et al., 2018b) and NWPU Campus (Cao et al., 2023).
ShanghaiTech (Liu et al., 2018b) breaks the single-scene limitations of UCSD Ped2 and CUHK Avenue, collecting 158 abnormal events from 13 scenes. Compared to the above datasets, it has a larger data scale and more complex anomaly types. Therefore, the performance of existing UVAD methods on ShanghaiTech is generally lower (AUC ), indicating the need for further exploration in cross-scene UVAD (Aich et al., 2023). Moreover, researchers proposed ShanghaiTech Weakly for WsVAD. Recent studies have shown that with similar scenarios and anomalies, the performance of WsVAD methods on the re-organized dataset significantly outperforms UVAD models on the original one. For example, mainstream WsVAD models like MIST (Feng et al., 2021a) and RTFM (Tian et al., 2021) achieve a frame-level AUC of up to 98%, surpassing state-of-the-art UVAD by nearly 20%. This phenomenon indicates that weakly supervised methods are more applicable in real-world scenarios.
Street Scene (Ramachandra and Jones, 2020) dataset is the first UVAD benchmark dedicated to anomaly detection in complex traffic, suitable for validating models’ deployment potential in smart city applications. The authors used static USB cameras to collect real street scenes (sidewalks, bike lanes, and roadways) in Cambridge, Massachusetts, USA, covering various traffic participants like pedestrians, non-motorized vehicles, and motor vehicles. Following the UVAD task setting, all 46 training videos depict normal events, while 35 test samples contain 205 abnormalities categorized into 17 classes. This dataset contains rich traffic behaviors, such as vehicle-centric driving, turning, and stopping, pedestrian-centric walking, jogging, and pushing strollers, and non-motorized vehicle-centric biking. Additionally, due to daylight angle changes, handling this dataset also requires addressing the effects of building shadows and moving backgrounds in urban scenes. Apart from temporal annotations, collectors also provided trajectory-level annotations, i.e., bounding boxes and IDs marking the boundaries of anomalies for fine-grained spatial localization learning.
NWPU Campus (Cao et al., 2023) is currently the largest UVAD dataset, containing 547 videos collected from 43 scenes, totaling 76.6GB. As shown in Table 3, this dataset’s data scale, scene quantity, and abnormal class diversity exceed those of existing unsupervised datasets, allowing for thorough testing of model effectiveness and generalization. It’s worth noting that the collectors considered the scene relevance of abnormal events, meaning similar behavioral patterns might be classified as different categories in different scenes. Therefore, during data collection, they simulated such behaviors extensively to enhance the dataset’s consistency with real-world scenes.
6.1.2. Weakly Supervised Datasets
UCF-Crime (Sultani et al., 2018) established the first benchmark for WsVAD and continuously promoted weakly supervised methods based on MIL. This dataset collected 1,900 unedited surveillance videos from the internet containing 850 crime events such as theft, robbery, and road accidents. Collectors categorized these anomalies into 13 classes, labeling videos containing such events as 1, while others were labeled as 0. This weak semantic annotation, termed video-level labels, is used to supervise model learning of finely detailed frame-level labels with solid semantics. Unlike UVAD, which only uses regular videos for model training, the UCF-Crime training set contains positive and negative samples, learning the differences between normal and abnormal by contrasting their pattern differences. We display abnormal samples from the UCF-Crime dataset in Fig. 8, including road accidents, shoplifting, stealing, shooting, and robbery. The anomalies defined in WsVAD align more with human recognition and have a higher degree of realism. Since predefined abnormal events are seen during the training phase, WsVAD exhibits more stable performance in detecting specific category anomalies, with less susceptibility to scene changes and external environmental influences. Subsequent weakly supervised datasets, such as XD-Violence (Wu et al., 2020) and these reconstructed from UVAD datasets like UCSD Ped2 and ShanghaiTech, also follow the data settings of UCF-Crime, with both training and testing sets containing positive and negative samples. However, the UCF-Crime dataset assumes VAD is a closed-set problem, as abnormal events in the test set have all appeared in the training set, thus unable to verify the model’s ability to discriminate anomalies in an open world.
XD-Violence (Wu et al., 2020) is the first mainstream large-scale VAD benchmark with multimodal data, including videos and synchronized audio. The authors collected various violent behaviors endangering personal safety from the internet as target anomalies, including riots, aggressions, shootings, and traffic accidents, as shown in Figure 8. Besides real surveillance videos from the internet, this dataset also includes audiovisual files from games and movies, significantly increasing the detection difficulty. XD-Violence opens a new avenue for VAD research, i.e., mining anomaly-related clues from heterogeneous data. Consistent with UCF-Crime (Sultani et al., 2018), XD-Violence provides coarse semantic video-level annotations. Each video only requires one label without annotating specific categories of anomalies, making the data preparation cost acceptable. Since the collectors did not consider the open-set characteristics of AD tasks, the anomaly event categories in the training and testing sets are consistent. Therefore, this dataset can only be used to test closed-set models. We recommend that future researchers place unseen anomalies in the testing set to develop multimodal VAD models with open-set detection capabilities.
6.1.3. Supervised Datasets
The high cost of collecting anomalies and fine-grained annotations makes existing VAD research typically avoid supervised learning requiring frame-level or pixel-level supervision. However, the rise of virtual data engines has made it feasible to construct diverse abnormal events and synchronously obtain precise labels. To this end, Acsintoae et al. (Acsintoae et al., 2022) constructed the first supervised VAD dataset named UBnormal by simulating virtual anomalies, providing a reference for addressing the data and label preparation challenges with virtual engines. In addition, this dataset can address the close-set concern of WsVAD tasks. The authors set up a validation set, including positive samples of different categories from the training set, to encourage models to detect unseen anomalies.
6.2. Evaluation Metrics
The existing works typically evaluate the proposed method from two perspectives: detection accuracy and model cost. On the one hand, models are expected to detect noteworthy anomalies as accurately as possible. Considering the varying influences of false positives and false negatives, along with the highly imbalanced data, researchers have proposed multiple quantitative metrics such as Area Under the Receiver Operating Characteristic curve (AUROC), Area Under Precision-Recall curve (AUPR), and Detection Rate, to assess model accuracy. Additionally, anomaly score curves are commonly used to qualitatively demonstrate the sensitivity to abnormal intervals, while prediction error maps are widely employed in reconstruction-based UVAD methods to visualize the performance of spatial localization. On the other hand, metrics like parameter size and inference speed determine whether the model can be deployed on resource-constrained devices. Thus, we categorize existing metrics based on the orientation and evaluation dimension, as illustrated in Fig. 9.
Accuracy-oriented metrics aim to evaluate a model’s ability to distinguish between normal and abnormal events, including quantitative metrics such as AUROC, AUPR, false alarm rate, and detection rate, as well as qualitative metrics like anomaly score curves and prediction error maps. While the task definitions vary across different VAD routes, they all aim to learn an anomaly detector capable of quantitatively measuring the abnormality level of test samples. Specifically, UVAD is a one-class classification task to train a model using regular events to describe normal patterns while considering all uncharacterizable samples as anomalies. The abnormality degree is computed by measuring the deviation to the learned normality model, which is typically normalized to the range of as anomaly scores. In contrast, WsVAD treats VAD as a regression task, using video-level labels to supervise fully connected networks directly outputting instance-level anomaly scores, similar to the FuVAD model. Therefore, despite differences in task settings and anomaly discrimination processes, metrics from binary classification tasks can be used to evaluate VAD models.
In most cases, the anomaly scores computed by NSVAD models are continuous values in the range , while the given data labels are binary discrete values, where 0 denotes negative (normal events) and 1 represents positive (anomalous events). Therefore, it’s necessary to select a threshold to convert relative abnormality scores into definitive binary labels for comparison. For example, with a threshold of 0.5, samples with scores lower than 0.5 are considered negative by the model, while those greater than or equal to 0.5 are considered positive. Thus, we can compute True Positive Rate (), False Positive Rate (), True Negative Rate (), and False Negative Rate (), as follows:
where , , , and represent correctly detected positive samples, negative samples misclassified as positive, correctly detected negative samples, and positive samples misclassified as negative, respectively.
Due to the highly imbalanced nature of positive and negative samples in NSVAD, some common evaluation metrics for classification tasks, such as accuracy, are not applicable. For example, a model biased towards outputting label 0 thus missing anomalous events, would be incorrectly evaluated as good under such metrics. Using a single threshold to simply assess a model’s ability to differentiate between normal and abnormal patterns is not wise. For instance, with a threshold of 0.5, a model that consistently outputs scores slightly below 0.5 for regular events and slightly above 0.5 for anomalies would be considered perfect because it would show optimal performance across various metrics. However, such a model may not have learned the inherent differences between normal and abnormal patterns well, resulting in minimal score gaps between the two, which could lead to failure in detecting subtle anomalies and normal instances with data bias in complex scenarios. Therefore, researchers have introduced Receiver Operating Characteristic (ROC) curves, which measure VAD models more comprehensively by selecting multiple thresholds. Specifically, this curve plots the True Negative Rate (TNR) against the True Positive Rate (TPR) at various thresholds. The Area Under the ROC curve (AUROC) has been the most widely used VAD evaluation metric. An ideal AUROC value of 1 indicates a model that outputs a score of 0 for all negative samples and 1 for all positive samples, aligning with our expectations. Considering that TN usually exceeds TP, researchers argue that Average Precision, i.e., the Area Under the Precision-Recall (AUPR) curve, is more suitable for evaluating anomaly detection tasks. The PR curve depicts precision and recall (i.e., TPR) at specific thresholds. The point on this curve where Precision equals Recall is the balance point. Currently, multi-modal anomaly detection models primarily use AP for quantitative evaluation. Anomaly score curve is commonly used to intuitively demonstrate the model’s response to anomalous events, presenting the temporal localization capability. In contrast, prediction error maps are often used to assess UVAD model’s spatial localization capability.
Current research primarily focuses on developing high-performance detection models while neglecting to evaluate the models’ deployment potential. A lightweight model is crucial in driving the application of NSVAD. Thus, we compile deployment-oriented metrics, including parameter size, Floating-Point Operations Per Second (FLOPS), and average inference speed. Specifically, parameter size indicates the number of learnable parameters, reflecting the complexity and storage cost of the model. In the real world, while complex models may offer performance gains, the resulting increase in memory and computational resource requirements may be unacceptable. Therefore, NSAVD should balance detection performance with model parameters. FLOPs represent the number of floating-point operations the model needs to perform during inference. This metric is crucial for end devices, as excessive FLOPs may lead to performance bottlenecks and reduced hardware lifespan. Some existing works report average inference speed, i.e., the number of frames the model can process during testing, to quantitatively reflect the model’s run time. However, due to inconsistent experimental environments, it cannot serve as an instructive and convincing metric in most cases. Due to space constraints and fairness of the comparison, we have collected the performance reported by existing methods but only organized these results in our GitHub11footnotemark: 1 repository for reference.
7. Research Cases
To intuitively demonstrate the deployment potential and challenges faced by NSVAD in real-world applications, we selected three of our research cases to apply NSVAD in modern industry (Liu et al., 2024b), smart cities (Liu et al., 2023c), and large-scale systems.
7.1. Appearance-Motion Prototype Network for Smart Industry
NSVAD can gather global spatial-temporal information in open spaces and detect various anomalous events in industrial settings to maintain production process safety and stability. Compared to traditional monitoring devices, video camera data offers stable quality and high information density, making NSVAD the preferred solution for Automatic Industrial Anomaly Detection Systems (AIADS). However, existing VAD research lacks in-depth exploration of industrial scenarios, and proposed methods often work only on outdoor public spaces, struggling to cope with the high-temporal dynamics and complex spatial-temporal interactions in manufacturing sites. On the one hand, the scarcity of industrial video datasets limits the development of AIADS. On the other hand, the diversity of industrial events and the complexity of patterns in such dynamic environments make developing efficient industrial NSVAD systems capable of effectively representing regular patterns and limitedly generalizing anomaly events a highly challenging task. To address this, inspired by the categorization of UVAD anomalies in Section 2.2 and motivated by the multi-stream approach introduced in Section 3.1., we propose in (Liu et al., 2024b) the first NSVAD model for AIADS, named the Appearance-Motion Prototype Network (AMP-Net), as shown in Fig. 10. Quantitative experiments on real-world industrial datasets demonstrate that AMP-Net can handle dynamic environments while maintaining limited generalization to various industrial anomalies.
As the first NSVAD case tailored for industrial scenarios, AMP-Net demonstrates the feasibility of using video for automated monitoring of production processes and the application prospects of NSVAD in the smart industry. The implementation of this model would not have been possible without previous explorations, particularly in multi-stream approaches (Liu et al., 2022c) and memory networks (Gong et al., 2019; Park et al., 2020), which also inspire future researchers to seek inspiration from existing solutions tailored to similar tasks when designing NSVAD systems for specific applications, as explained in Section 3.1 of Supplement. Although the model achieves excellent performance on self-collected industrial datasets, it fails to consider the deployment cost on terminal devices and the arbitrary data shifts in real-world environments, indicating that there is still a distance to go before it can be truly applied in modern factories. Additionally, quantitatively analyzing the model’s detection capabilities for different industrial anomaly events to develop interpretable NSVAD systems and improve the model’s detection capabilities for highly hazardous anomalies is also a realistic requirement worth exploring.
7.2. Causal Video Normality Learning Towards Smart Cities
Smart cities are reshaping human lifestyles, relying on data sensing and understanding technicals driven by AI and next-generation communication. Among the vast and multimodal data generated by typical smart city applications such as modern transportation, smart homes, and public services, videos often dominate data volume and information density, making automated video processing a fundamental task in smart city research. NSVAD, tailored for smart cities, holds the promise of providing new solutions for traffic accident detection, elderly state monitoring, and public safety protection. However, challenges such as label-independent data shifts and subtle anomalies brought about by the diversity of events in such complex scenarios need to be addressed. Despite efforts by UVAD methods (Liu et al., 2018b; Gong et al., 2019; Chang et al., 2020), including the AMP-Net (Liu et al., 2024b), to balance the representation and generalization capabilities, deep learning-based models have struggled to establish simple mapping relationships between regular events and labels, lacking exploration of the causal relationships in video regularity. Our case studies in (Liu et al., 2023c) indicate that deep representation learning (DRL) struggles to learn the essential differences between normal and abnormal events and fails to resist arbitrary biases in smart cities, leading to significant performance degradation when handling complex regular events and minor anomalies.
To this end, we introduced Causal Representation Learning (CRL) into NSVAD, proposing Causal Video Normality Learning (CVNL) to explore potential causal relationships (Liu et al., 2023c). Compared to the shallow description of video normality based on statistical dependency established by DRL, CVNL attempts to discern anomalies based on the inherent causal consistency of regular events, enabling NSVAD to generalize reasonably well to diverse regular events with label-independent biases while remaining sensitive to anomalies. Extensive experiments on large-scale cross-scene datasets demonstrate that CVNL can meet the detection requirements in complex environments like smart cities. Even when only a few regular events are available, it can maintain stable performance. The simplified framework of CVNL is outlined in Fig. 11, and more information is provided in the paper (Liu et al., 2023c) and Section 3.2 of Supplement.
7.3. End-to-Cloud Collaborative Framework for NSVAD
Inspired by NSAI (Song et al., 2022), we delineate the research on deployable NSVAD systems for real-world, large-scale IoT and mobile internet applications into a hierarchical structure as shown in Fig. 12, comprising hardware layer, system layer, algorithm layer, and application layer. Existing NSVAD works mainly focus on the algorithm and application, emphasizing model development for specific scenarios and aiming to design customized solutions such as unsupervised crowd anomaly discrimination, weakly supervised crime identification, and multimodal violence detection. Due to the limited computational resources of terminal cameras, such models typically rely on edge devices of local monitoring systems to perform data analysis and anomaly detection, resulting in limited spatial scale and temporal length of detectable anomalies. Taking traffic management in smart cities as an example, constructing NSVAD systems for millions of traffic cameras across the entire metropolitan area is necessary for detecting traffic congestion affecting overall road conditions and long-term abnormal traffic flows. However, existing methods can only achieve local short-term detection at the intersection level or even on single terminal devices. Although the development of mobile communication technology enables smart terminals such as mobile phones and tablets to utilize cloud servers for device-interconnected large-scale model learning and global detection, existing cloud-based mobile applications typically upload raw data, which is feasible for NSVAD. On the one hand, the information density of video data and the data set that can be collected per unit of time are higher than other modalities, leading to a significant increase in communication overhead for mobile devices. On the other hand, anomaly detection tasks often involve sensitive information, and directly transmitting such data to the cloud may raise concerns about user privacy among the public. To address this, to balance the communication overhead between terminal devices and cloud servers and eliminate data security risks, we design a configurable end-to-cloud collaborative framework.
Several end devices and a cloud server collaborate to accomplish various types of anomaly event detection, including local, global, instantaneous, and long-term, for large-scale terminal systems. Specifically, monitoring cameras or mobile devices on the end-side can serve as computing relays for cloud servers to perform preprocessing and feature extraction tasks. Thus, end devices tailored for specific scenarios or users only need to transmit data features or intermediate results to the cloud, significantly reducing sensitivity compared to raw visual information. The cloud server aggregates features and results from various edn devices and constructs large models (Liu et al., 2023b) with global and long-term detection capabilities, unloading models to local devices or directly reporting results to the end according to scene requirements. Our end-cloud collaborative NSVAD architecture is flexible and configurable, supporting an indefinite number of devices to join, and the execution party of each task can freely choose between end and cloud according to actual needs (such as real-time performance and complexity). Thus, this architecture can leverage the natural advantages of devices close to raw data, thereby reducing latency and communication overhead, while also utilizing cloud computing power to build high-performance large models without raising concerns about data privacy.
The core modules of the ECC-NSVAD system include a deployment platform, data pipeline, and computing containers. Computing containers aim to provide a cross-platform and lightweight execution environment for various NSVAD models, supporting continuous updates of models. The deployment platform directly targets terminal devices such as monitoring cameras and smartphones, deploying large NSVAD models learned from the cloud locally, similar to existing end-cloud collaborative architectures used for e-commerce recommendation services. The data pipeline is based on the Flink stream processing framework, which directly uploads all raw data collected by devices to the cloud, leading to high redundancy, large communication overhead, and high privacy sharing. To address this, we propose a lightweight data pipeline referring to Alibaba’s Walle framework, which optimizes resource overhead through end-cloud collaboration at the data level. The real-time channel adopts long-link communication and optimizes SSL protocol to reduce latency. In practical applications, ECC-NSVAD utilizes lightweight models to execute instantaneous local anomaly behavior detection on end devices and transfers features of low-confidence segments to the cloud for detecting potential complex anomalies. Case studies show that compared to traditional data center solutions (where all videos are directly transmitted to the cloud for central model processing), our end-cloud collaborative framework reduces cloud workload by 78%, increases detection accuracy (frame-level AUC) by 12%, and reduces false positive rates by 46% in complex dynamic environments.
8. Discussion and Summary
8.1. Research Challenges
In Sections 3-5, we have briefly outlined scientific problems addressed by various algorithms when presenting their motivation and implementation. To further help readers grasp the key challenges in NSVAD research, we detailed the unresolved and impending problems from the perspectives of data, labels, models, and systems. Unlike previous works focusing solely on algorithm design, we also discussed the cost bottlenecks that NSVAD deployments may encounter in real-world applications, aiming to guide researchers from different fields.
8.1.1. Data
Due to variations in scenes, camera view, and device configurations, real-world videos often exhibit label-independent domain shifts, where videos with the same semantic labels manifest slight differences in spatial-temporal patterns. Although such variations are typically easily understood by humans, they may serve as the basis for discerning anomalies by NSVAD models, leading to high false positive rates. Existing methods often validate models on datasets collected from a single scene with the same device and view to circumvent this issue. Although the UCSD dataset (Li et al., 2013) captures videos from two perspectives, parallel to and perpendicular to intersections, it is divided into two datasets for learning view-specific NSVAD models. Additionally, the ShanghaiTech dataset (Liu et al., 2018b) comprises 13 scenes, yet existing methods often overlook scene differences, treating it as single-scene, resulting in significantly lower performance compared to simple datasets like UCSD Ped2 (Li et al., 2013) and CUHK Avenue (Lu et al., 2013). In the real world, variations in scenes and views, as well as disparities between surveillance cameras and smartphones, are inevitable, making it unrealistic to learn specialized models for each specific setup. Given the rarity of anomalies, recent researchers have proposed using virtual data engines to simulate anomaly events to construct rich positive sample sets. Additionally, the XD-Violence dataset (Wu et al., 2020) introduces movie and game scenes. While such data enables the training of large-scale models, they often differ from real anomalies, rendering trained models ineffective in real-world applications. Therefore, bridging the domain shift between virtual anomalies and the real world is crucial for deployable NSVAD models. Despite becoming a hot topic in anomaly detection and multimodal understanding communities, multimodal NSVAD is still limited to the fusion of RGB images and synchronized audio. Existing methods focus solely on traditional image, audio, and optical flow modalities, unable to extract anomaly clues from new modalities like language and subtitle text, limiting NSVAD applications in streaming media and video broadcasting.
8.1.2. Label
Despite the repeated emphasis in the anomaly detection community on the rarity and diversity of anomaly samples, mainstream anomaly detection models often train in an unsupervised manner using only normal samples. However, as stated in Section 2.1, unsupervised methods actually require all training sample labels to be 0. Therefore, challenges in the label aspect include the impact of noisy labels on model performance and the annotation cost of large-scale datasets. Specifically, UVAD relies on additional data filtering to avoid mixing anomaly samples in the training set, making it unable to perform online learning directly from raw video streams like FuVAD and continuously improve model detection capabilities in real-world environments. Another type, WsVAD, relies on video-level labels to avoid high labor costs. Future work should focus on improving the stability of FuVAD and its performance in complex environments or developing novel hybrid supervised models to eliminate the influence of label noise.
8.1.3. Model
Unsupervised NSVAD benefits from efficient representation of spatial-temporal features by video self-supervised learning models but may also miss anomalies due to over-generalization. Therefore, the key of unsupervised methods lies in balancing representation and generalization to simultaneously reduce false positives and false negatives. Cases in Fig. 10 and 11 demonstrate the feasibility of memory networks and causal representation learning to achieve this goal, but performance differences on complex datasets indicate that improving the adaptability of UVAD in real environments remains a challenge. Additionally, WsVAD can only detect predefined specific anomaly events, violating the open-world assumption. This not only limits the application scenarios of the model but also hinders the collaborative development of NSVAD technology, leading existing reviews to typically not consider it within the scope of VAD research. Although existing research indicates that the detection results of WsVAD are more reliable compared to UVAD and FuVAD, its core ranking loss is not convincingly accepted. To further promote VAD technology in smart surveillance systems and mobile internet applications, it is necessary to develop interpretable models and rely on NSVAD systems with data encryption capabilities.
8.1.4. System
Existing NSVAD research focuses solely on improving performance on existing datasets while ignoring a series of challenges in model deployment, such as the compatibility of heterogeneous devices and resource costs. Specifically, NSVAD for different businesses (such as surveillance video analysis, movie content monitoring, and non-real video identification) needs to deal with different task settings and data modalities. However, existing work mainly focuses on anomaly detection in surveillance videos while neglecting new applications like short videos and live streaming. Additionally, terminal devices, especially mobile phones, often have limited storage, computing, and communication resources, so NSVAD model design needs to move towards lightweight and shift from algorithm to system to achieve synchronous optimization of detection performance and cost.
8.2. Trends and Opportunities
Based on the deployable NSVAD partitioned architecture depicted in Fig. 12, we have summarized the trends and open opportunities from hardware, system, algorithm, and application layers by incorporating the development trends of artificial intelligence and communication technologies, as well as the application requirements of NSVAD in smart cities and mobile internet, aiming to guide researchers in various fields to engage in relevant work.
8.2.1. Hardware Layer
The deployment of billions of cameras in roads, factories, and public places not only provides diverse application scenarios but also offers ample data support for training large-scale NSVAD models. Next-generation communication devices and image processing units can provide communication and computational support for building large-scale NSVAD systems, making it feasible for cloud-based global model learning in end-cloud collaborative architecture. Additionally, emerging sensing devices such as thermal imaging, motion cameras, and event cameras will expand the application potential of NSVAD in scenarios like military and sports. Therefore, we believe that the advancement and innovation of hardware devices strongly drive the development and application of NSVAD technology, and developing deployable NSVAD systems for new types of sensors and large-scale IoT systems will become a trend.
8.2.2. System Layer
The system layer aims to bridge the hardware and algorithm layers, providing interfaces for model deployment on terminal devices and supporting the collaborative optimization of computation and communication execution in NSVAD systems. Previous NSVAD research focused on algorithm design while neglecting system layer development for a long time. The end-cloud collaborative framework introduced in Section 7.3 explicitly indicates the importance of the system layer in promoting the video anomaly detection technology in large-scale IoT and mobile internet applications and designs the first NSVAD system layer for end-cloud collaborative detection and resource allocation. With the further development of edge artificial intelligence and mobile communication networks, new solutions will be sought for communication strategy optimization, computation offloading, distributed model aggregation, and interface flexibility faced by the system layer, promoting NSVAD towards integration and intelligence.
8.2.3. Algorithm Layer
Combining the latest advancements and domain concerns, we believe the development opportunities for algorithms include: 1) the introduction of large-scale heterogeneous datasets; 2) the adaptability transfer of efficient representation learning methods; 3) innovative combinations of emerging artificial intelligence tasks with NSVAD; and 4) assistance from implicit knowledge in large models (Han et al., 2024a). Specifically, expanding business scenarios from smart transportation and modern factories provide rich data sources for VSVAD, making it possible to train large-scale models. The rise of online video networks and streaming platforms provides additional data modalities beyond images, such as audio, subtitles, and language, aiding in exploring complex anomaly clues. Furthermore, the development of virtual data engines allows easy emulation of rare anomaly events and provides pixel-level annotations. The rarity and diversity of anomalies and the difficulty of collecting cross-scene videos will no longer be bottlenecks for NSVAD model development. Rich data modalities and volume will drive the development of multimodal and cross-scene NSVAD models and are expected to foster supervised interpretable research routes. The development of deep learning will free NSVAD algorithms from manual features and enable them to extract spatial-temporal features end-to-end and model video normality, while the continuous advancement of deep neural networks (such as attention, Transformer, masked autoencoders) and emerging representation learning methods (such as video self-supervised learning and causal representation learning) will further improve NSVAD’s learning capability. The advancement of emerging artificial intelligence technologies will provide references for addressing specific concerns of NSVAD. For example, unsupervised NSVAD recovers from label-agnostic data biases in complex environments, often encountering significant performance degradation when dealing with diverse common events, a problem long studied in domain generalization tasks. We believe the combination of domain generalization and NSVAD research will help address the negative impact of data bias in unsupervised solutions. Data security and privacy protection have always been concerns for users and researchers when building deployable NSVAD systems for large-scale IoT. Federated learning will provide feasible solutions. The rise of large models will propel generative artificial intelligence research to a climax and give birth to numerous phenomenon-level applications, reshaping industries, and NSVAD is no exception. We believe that large models contain implicit knowledge related to anomalous events, which is crucial for understanding the fundamental differences between normal and abnormal and developing interpretable NSVAD models.
8.2.4. Application Layer
On the one hand, applications such as smart transportation and live content monitoring provide ample validation scenarios for NSVAD models, driving the development of cross-scene perspective robust models for real-world deployment. On the other hand, business requirements in specific scenarios will spur new research tasks. For example, thermal imaging, as a completely passive sensing method, has been widely used in autonomous driving and the modern military to overcome the limitations of optical cameras at night. NSVAD model design based on thermal sensing devices will address texture loss and boundary entanglement caused by blackbody radiation, facilitating all-weather anomaly event detection. Additionally, the limited resources of terminal devices and the perceptual range of local systems indicate that lightweight models and the development of end-cloud collaborative NSVAD systems for large-scale applications are worth exploring.
8.3. Summary
NSVAD has emerged as a significant area of research with broad implications for smart cities and the mobile internet, exerting essential influence across various domains such as traffic management, industrial manufacturing, and the operations of online video platforms. This influence is critical for maintaining urban safety and ensuring a clear cyberspace. Originating from the confluence of anomaly detection and video understanding, NSVAD has expanded beyond mere algorithm design, transforming into a multifaceted subject of interest that spans AI, IoT, and computing. As the pioneering tutorial-type paper on NSVAD, this article comprehensively outlines its research landscape, clarifying the foundational concepts and developmental trajectories across various research avenues. In particular, we examine recent advancements in unsupervised, weakly supervised, and fully unsupervised methods, providing detailed explanations of classical solutions. Remarkably, this article centrally presents our latest explorations to NSVAD in modern industry, smart cities, and complex systems. Finally, leveraging our experiences, we analyze the challenges, trends, and opportunities within the future vision of NSVAD, aiming to spark inspiration for following engineers and researchers.
Acknowledgements.
This work has been supported in part by the China Mobile Research Fund of the Chinese Ministry of Education under Grant KEH2310029, the National Natural Science Foundation of China under Grant 62250410368, and the Specific Research Fund of the Innovation Platform for Academicians of Hainan Province under Grant YSPTZX202314. Additionally, this work has received backing from the Shanghai Key Research Laboratory of NSAI and the Joint Laboratory on Networked AI Edge Computing Fudan University-Changan. The work of Yang Liu is financially supported in part by the China Scholarship Council (File No. 202306100221). The authors would like to express their gratitude to the anonymous reviewers for their insightful comments and valuable suggestions. We appreciate Hao Yang, Liangyu Teng, and Yuntian Shi for their invaluable assistance and thorough review of this article.References
- (1)
- Acsintoae et al. (2022) Andra Acsintoae, Andrei Florescu, Mariana-Iuliana Georgescu, Tudor Mare, Paul Sumedrea, Radu Tudor Ionescu, Fahad Shahbaz Khan, and Mubarak Shah. 2022. UBnormal: New Benchmark for Supervised Open-Set Video Anomaly Detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 20143–20153.
- Adam et al. (2008) Amit Adam, Ehud Rivlin, Ilan Shimshoni, and Daviv Reinitz. 2008. Robust real-time unusual event detection using multiple fixed-location monitors. IEEE transactions on pattern analysis and machine intelligence 30, 3 (2008), 555–560.
- Aich et al. (2023) Abhishek Aich, Kuan-Chuan Peng, and Amit K Roy-Chowdhury. 2023. Cross-domain video anomaly detection without target domain adaptation. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 2579–2591.
- Bao et al. (2022) Qianyue Bao, Fang Liu, Yang Liu, Licheng Jiao, Xu Liu, and Lingling Li. 2022. Hierarchical scene normality-binding modeling for anomaly detection in surveillance videos. In Proceedings of the 30th ACM International Conference on Multimedia. 6103–6112.
- Basak and Gautam (2024) Suvramalya Basak and Anjali Gautam. 2024. Diffusion-based normality pre-training for weakly supervised video anomaly detection. Expert Systems with Applications (2024), 124013.
- Benezeth et al. (2009) Yannick Benezeth, P-M Jodoin, Venkatesh Saligrama, and Christophe Rosenberger. 2009. Abnormal events detection based on spatio-temporal co-occurences. In 2009 IEEE conference on computer vision and pattern recognition. IEEE, 2458–2465.
- Blázquez-García et al. (2021) Ane Blázquez-García, Angel Conde, Usue Mori, and Jose A Lozano. 2021. A review on outlier/anomaly detection in time series data. ACM Computing Surveys (CSUR) 54, 3 (2021), 1–33.
- Boukerche et al. (2020) Azzedine Boukerche, Lining Zheng, and Omar Alfandi. 2020. Outlier detection: Methods, models, and classification. ACM Computing Surveys (CSUR) 53, 3 (2020), 1–37.
- Cai et al. (2021b) Ruichu Cai, Hao Zhang, Wen Liu, Shenghua Gao, and Zhifeng Hao. 2021b. Appearance-motion memory consistency network for video anomaly detection. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 35. 938–946.
- Cai et al. (2021a) Yiheng Cai, Jiaqi Liu, Yajun Guo, Shaobin Hu, and Shinan Lang. 2021a. Video anomaly detection with multi-scale feature and temporal information fusion. Neurocomputing 423 (2021), 264–273.
- Cao et al. (2023) Congqi Cao, Yue Lu, Peng Wang, and Yanning Zhang. 2023. A New Comprehensive Benchmark for Semi-supervised Video Anomaly Detection and Anticipation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 20392–20401.
- Cao et al. (2024) Congqi Cao, Yue Lu, and Yanning Zhang. 2024. Context recovery and knowledge retrieval: A novel two-stream framework for video anomaly detection. IEEE Transactions on Image Processing (2024).
- Carreira and Zisserman (2017) Joao Carreira and Andrew Zisserman. 2017. Quo vadis, action recognition? a new model and the kinetics dataset. In proceedings of the IEEE Conference on Computer Vision and Pattern Recognition. 6299–6308.
- Chandrakala et al. (2022) S Chandrakala, K Deepak, and G Revathy. 2022. Anomaly detection in surveillance videos: a thematic taxonomy of deep models, review and performance analysis. Artificial Intelligence Review (2022), 1–50.
- Chang et al. (2021) Yunpeng Chang, Zhigang Tu, Wei Xie, Bin Luo, Shifu Zhang, Haigang Sui, and Junsong Yuan. 2021. Video anomaly detection with spatio-temporal dissociation. Pattern Recognition 122 (2021), 108213.
- Chang et al. (2020) Yunpeng Chang, Zhigang Tu, Wei Xie, and Junsong Yuan. 2020. Clustering driven deep autoencoder for video anomaly detection. In European Conference on Computer Vision. Springer, 329–345.
- Chen et al. (2022) Chengwei Chen, Yuan Xie, Shaohui Lin, Angela Yao, Guannan Jiang, Wei Zhang, Yanyun Qu, Ruizhi Qiao, Bo Ren, and Lizhuang Ma. 2022. Comprehensive Regularization in a Bi-directional Predictive Network for Video Anomaly Detection. In Proceedings of the American association for artificial intelligence. 1–9.
- Chen (2020) Chang Wen Chen. 2020. Internet of video things: Next-generation IoT with visual sensors. IEEE Internet of Things Journal 7, 8 (2020), 6676–6685.
- Chen et al. (2020) Dongyue Chen, Pengtao Wang, Lingyi Yue, Yuxin Zhang, and Tong Jia. 2020. Anomaly detection in surveillance video based on bidirectional prediction. Image and Vision Computing 98 (2020), 103915.
- Chen et al. (2021) Dongyue Chen, Lingyi Yue, Xingya Chang, Ming Xu, and Tong Jia. 2021. NM-GAN: Noise-modulated generative adversarial network for video anomaly detection. Pattern Recognition 116 (2021), 107969.
- Cheng et al. (2023) Kai Cheng, Yang Liu, and Xinhua Zeng. 2023. Learning graph enhanced spatial-temporal coherence for video anomaly detection. IEEE Signal Processing Letters 30 (2023), 314–318.
- Cheng et al. (2015) Kai-Wen Cheng, Yie-Tarng Chen, and Wen-Hsien Fang. 2015. Video anomaly detection and localization using hierarchical feature representation and Gaussian process regression. In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition. 2909–2917.
- Chong and Tay (2017) Yong Shean Chong and Yong Haur Tay. 2017. Abnormal event detection in videos using spatiotemporal autoencoder. In International symposium on neural networks. Springer, 189–196.
- Christiansen et al. (2016) Peter Christiansen, Lars N Nielsen, Kim A Steen, Rasmus N Jørgensen, and Henrik Karstoft. 2016. DeepAnomaly: Combining background subtraction and deep learning for detecting obstacles and anomalies in an agricultural field. Sensors 16, 11 (2016), 1904.
- Cook et al. (2019) Andrew A Cook, Göksel Mısırlı, and Zhong Fan. 2019. Anomaly detection for IoT time-series data: A survey. IEEE Internet of Things Journal 7, 7 (2019), 6481–6494.
- Cui et al. (2011) Xinyi Cui, Qingshan Liu, Mingchen Gao, and Dimitris N Metaxas. 2011. Abnormal detection using interaction energy potentials. In CVPR 2011. IEEE, 3161–3167.
- Dalal and Triggs (2005) Navneet Dalal and Bill Triggs. 2005. Histograms of oriented gradients for human detection. In 2005 IEEE computer society conference on computer vision and pattern recognition (CVPR’05), Vol. 1. Ieee, 886–893.
- Dalal et al. (2006) Navneet Dalal, Bill Triggs, and Cordelia Schmid. 2006. Human detection using oriented histograms of flow and appearance. In European conference on computer vision. Springer, 428–441.
- Deepak et al. (2021) K Deepak, S Chandrakala, and C Krishna Mohan. 2021. Residual spatiotemporal autoencoder for unsupervised video anomaly detection. Signal, Image and Video Processing 15, 1 (2021), 215–222.
- Dollár et al. (2005) Piotr Dollár, Vincent Rabaud, Garrison Cottrell, and Serge Belongie. 2005. Behavior recognition via sparse spatio-temporal features. In 2005 IEEE international workshop on visual surveillance and performance evaluation of tracking and surveillance. IEEE, 65–72.
- Dong et al. (2020) Fei Dong, Yu Zhang, and Xiushan Nie. 2020. Dual discriminator generative adversarial network for video anomaly detection. IEEE Access 8 (2020), 88170–88176.
- Doshi and Yilmaz (2021) Keval Doshi and Yasin Yilmaz. 2021. Online anomaly detection in surveillance videos with asymptotic bound on false alarm rate. Pattern Recognition 114 (2021), 107865.
- Ektefaie et al. (2023) Yasha Ektefaie, George Dasoulas, Ayush Noori, Maha Farhat, and Marinka Zitnik. 2023. Multimodal learning with graphs. Nature Machine Intelligence 5, 4 (2023), 340–350.
- Fan et al. (2020) Yaxiang Fan, Gongjian Wen, Deren Li, Shaohua Qiu, Martin D Levine, and Fei Xiao. 2020. Video anomaly detection and localization via gaussian mixture fully convolutional variational autoencoder. Computer Vision and Image Understanding 195 (2020), 102920.
- Feng et al. (2021a) Jia-Chang Feng, Fa-Ting Hong, and Wei-Shi Zheng. 2021a. Mist: Multiple instance self-training framework for video anomaly detection. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 14009–14018.
- Feng et al. (2021b) Xinyang Feng, Dongjin Song, Yuncong Chen, Zhengzhang Chen, Jingchao Ni, and Haifeng Chen. 2021b. Convolutional transformer based dual discriminator generative adversarial networks for video anomaly detection. In Proceedings of the 29th ACM International Conference on Multimedia. 5546–5554.
- Geng et al. (2020) Chuanxing Geng, Sheng-jun Huang, and Songcan Chen. 2020. Recent advances in open set recognition: A survey. IEEE transactions on pattern analysis and machine intelligence 43, 10 (2020), 3614–3631.
- Georgescu et al. (2021a) Mariana-Iuliana Georgescu, Antonio Barbalau, Radu Tudor Ionescu, Fahad Shahbaz Khan, Marius Popescu, and Mubarak Shah. 2021a. Anomaly detection in video via self-supervised and multi-task learning. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 12742–12752.
- Georgescu et al. (2021b) Mariana Iuliana Georgescu, Radu Tudor Ionescu, Fahad Shahbaz Khan, Marius Popescu, and Mubarak Shah. 2021b. A background-agnostic framework with adversarial training for abnormal event detection in video. IEEE transactions on pattern analysis and machine intelligence 44, 9 (2021), 4505–4523.
- Girshick (2015) Ross Girshick. 2015. Fast r-cnn. In Proceedings of the IEEE international conference on computer vision. 1440–1448.
- Gong et al. (2019) Dong Gong, Lingqiao Liu, Vuong Le, Budhaditya Saha, Moussa Reda Mansour, Svetha Venkatesh, and Anton van den Hengel. 2019. Memorizing normality to detect anomaly: Memory-augmented deep autoencoder for unsupervised anomaly detection. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 1705–1714.
- Han et al. (2024b) Xingshuo Han, Xiao Wang, Kui Jiang, Wei Liu, Ruimin Hu, Xuefeng Pan, and Xin Xu. 2024b. Mutuality Attribute Makes Better Video Anomaly Detection. In ICASSP 2024-2024 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 2670–2674.
- Han et al. (2024a) Zeyu Han, Chao Gao, Jinyang Liu, Sai Qian Zhang, et al. 2024a. Parameter-Efficient Fine-Tuning for Large Models: A Comprehensive Survey. arXiv preprint arXiv:2403.14608 (2024).
- Hasan et al. (2016) Mahmudul Hasan, Jonghyun Choi, Jan Neumann, Amit K Roy-Chowdhury, and Larry S Davis. 2016. Learning temporal regularity in video sequences. In Proceedings of the IEEE conference on computer vision and pattern recognition. 733–742.
- He et al. (2022) Kaiming He, Xinlei Chen, Saining Xie, Yanghao Li, Piotr Dollár, and Ross Girshick. 2022. Masked autoencoders are scalable vision learners. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 16000–16009.
- Herrera et al. (2016) Francisco Herrera, Sebastián Ventura, Rafael Bello, Chris Cornelis, Amelia Zafra, Dánel Sánchez-Tarragó, Sarah Vluymans, Francisco Herrera, Sebastián Ventura, Rafael Bello, et al. 2016. Multiple instance learning. Springer.
- Hinami et al. (2017) Ryota Hinami, Tao Mei, and Shin’ichi Satoh. 2017. Joint detection and recounting of abnormal events by learning deep generic knowledge. In Proceedings of the IEEE international conference on computer vision. 3619–3627.
- Hojjati et al. (2024) Hadi Hojjati, Thi Kieu Khanh Ho, and Naregs Armanfard. 2024. Self-supervised anomaly detection in computer vision and beyond: A survey and outlook. Neural Networks (2024), 106106.
- Hu et al. (2022) Jingtao Hu, Guang Yu, Siqi Wang, En Zhu, Zhiping Cai, and Xinzhong Zhu. 2022. Detecting Anomalous Events from Unlabeled Videos via Temporal Masked Auto-Encoding. In 2022 IEEE International Conference on Multimedia and Expo (ICME). IEEE, 1–6.
- Hu et al. (2016) Xing Hu, Shiqiang Hu, Yingping Huang, Huanlong Zhang, and Hanbing Wu. 2016. Video anomaly detection using deep incremental slow feature analysis network. IET Computer Vision 10, 4 (2016), 258–267.
- Hu et al. (2018) Xing Hu, Yingping Huang, Xiumin Gao, Lingkun Luo, and Qianqian Duan. 2018. Squirrel-cage local binary pattern and its application in video anomaly detection. IEEE Transactions on Information Forensics and Security 14, 4 (2018), 1007–1022.
- Huang et al. (2021) Chao Huang, Zhihao Wu, Jie Wen, Yong Xu, Qiuping Jiang, and Yaowei Wang. 2021. Abnormal event detection using deep contrastive learning for intelligent video surveillance system. IEEE Transactions on Industrial Informatics 18, 8 (2021), 5171–5179.
- Ionescu et al. (2019) Radu Tudor Ionescu, Fahad Shahbaz Khan, Mariana-Iuliana Georgescu, and Ling Shao. 2019. Object-centric auto-encoders and dummy anomalies for abnormal event detection in video. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 7842–7851.
- Jedari et al. (2020) Behrouz Jedari, Gopika Premsankar, Gazi Illahi, Mario Di Francesco, Abbas Mehrabi, and Antti Ylä-Jääski. 2020. Video caching, analytics, and delivery at the wireless edge: A survey and future directions. IEEE Communications Surveys & Tutorials 23, 1 (2020), 431–471.
- Jiang et al. (2021) Xiantao Jiang, F Richard Yu, Tian Song, and Victor CM Leung. 2021. A survey on multi-access edge computing applied to video streaming: Some research issues and challenges. IEEE Communications Surveys & Tutorials 23, 2 (2021), 871–903.
- Jing and Tian (2020) Longlong Jing and Yingli Tian. 2020. Self-supervised visual feature learning with deep neural networks: A survey. IEEE transactions on pattern analysis and machine intelligence 43, 11 (2020), 4037–4058.
- Kiran et al. (2018) B Ravi Kiran, Dilip Mathew Thomas, and Ranjith Parakkal. 2018. An overview of deep learning based methods for unsupervised and semi-supervised anomaly detection in videos. Journal of Imaging 4, 2 (2018), 36.
- Ko and Sim (2018) Kwang-Eun Ko and Kwee-Bo Sim. 2018. Deep convolutional framework for abnormal behavior detection in a smart surveillance system. Engineering Applications of Artificial Intelligence 67 (2018), 226–234.
- Kong and Fu (2022) Yu Kong and Yun Fu. 2022. Human action recognition and prediction: A survey. International Journal of Computer Vision 130, 5 (2022), 1366–1401.
- Li et al. (2024) Chaobo Li, Hongjun Li, and Guoan Zhang. 2024. Cross-modality integration framework with prediction, perception and discrimination for video anomaly detection. Neural Networks 172 (2024), 106138.
- Li et al. (2022) Daoheng Li, Xiushan Nie, Xiaofeng Li, Yu Zhang, and Yilong Yin. 2022. Context-related video anomaly detection via generative adversarial network. Pattern Recognition Letters 156 (2022), 183–189.
- Li et al. (2020) Nanjun Li, Faliang Chang, and Chunsheng Liu. 2020. Spatial-temporal cascade autoencoder for video anomaly detection in crowded scenes. IEEE Transactions on Multimedia 23 (2020), 203–215.
- Li et al. (2021a) Tong Li, Xinyue Chen, Fushun Zhu, Zhengyu Zhang, and Hua Yan. 2021a. Two-stream deep spatial-temporal auto-encoder for surveillance video abnormal event detection. Neurocomputing 439 (2021), 256–270.
- Li et al. (2021c) Tangqing Li, Zheng Wang, Siying Liu, and Wen-Yan Lin. 2021c. Deep unsupervised anomaly detection. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 3636–3645.
- Li et al. (2013) Weixin Li, Vijay Mahadevan, and Nuno Vasconcelos. 2013. Anomaly detection and localization in crowded scenes. IEEE transactions on pattern analysis and machine intelligence 36, 1 (2013), 18–32.
- Li et al. (2019) Yuanyuan Li, Yiheng Cai, Jiaqi Liu, Shinan Lang, and Xinfeng Zhang. 2019. Spatio-Temporal Unity Networking for Video Anomaly Detection. IEEE Access 7 (2019), 172425–172432.
- Li et al. (2021b) Zewen Li, Fan Liu, Wenjie Yang, Shouheng Peng, and Jun Zhou. 2021b. A survey of convolutional neural networks: analysis, applications, and prospects. IEEE transactions on neural networks and learning systems (2021).
- Li et al. (2023) Zhong Li, Yuxuan Zhu, and Matthijs Van Leeuwen. 2023. A survey on explainable anomaly detection. ACM Transactions on Knowledge Discovery from Data 18, 1 (2023), 1–54.
- Liu et al. (2018b) Wen Liu, Weixin Luo, Dongze Lian, and Shenghua Gao. 2018b. Future frame prediction for anomaly detection–a new baseline. In Proceedings of the IEEE conference on computer vision and pattern recognition. 6536–6545.
- Liu et al. (2023a) Yang Liu, Zhengliang Guo, Jing Liu, Chengfang Li, and Liang Song. 2023a. Osin: Object-centric scene inference network for unsupervised video anomaly detection. IEEE Signal Processing Letters 30 (2023), 359–363.
- Liu et al. (2024a) Yang Liu, Bobo Ju, Dingkang Yang, Liyuan Peng, Di Li, Peng Sun, Chengfang Li, Hao Yang, Jing Liu, and Liang Song. 2024a. Memory-enhanced spatial-temporal encoding framework for industrial anomaly detection system. Expert Systems with Applications (2024), 123718. https://doi.org/10.1016/j.eswa.2024.123718
- Liu et al. (2018a) Yusha Liu, Chun-Liang Li, and Barnabás Póczos. 2018a. Classifier Two Sample Test for Video Anomaly Detections.. In BMVC. 71.
- Liu et al. (2021a) Yang Liu, Shuang Li, Jing Liu, Hao Yang, Mengyang Zhao, Xinhua Zeng, Wei Ni, and Liang Song. 2021a. Learning Attention Augmented Spatial-temporal Normality for Video Anomaly Detection. In 2021 3rd International Symposium on Smart and Healthy Cities (ISHC). IEEE, 137–144.
- Liu et al. (2022a) Yang Liu, Jing Liu, Jieyu Lin, Mengyang Zhao, and Liang Song. 2022a. Appearance-Motion United Auto-Encoder Framework for Video Anomaly Detection. IEEE Transactions on Circuits and Systems II: Express Briefs 69, 5 (2022), 2498–2502.
- Liu et al. (2022b) Yang Liu, Jing Liu, Wei Ni, and Liang Song. 2022b. Abnormal Event Detection with Self-guiding Multi-instance Ranking Framework. In 2022 International Joint Conference on Neural Networks (IJCNN). IEEE, 01–07.
- Liu et al. (2024b) Yang Liu, Jing Liu, Kun Yang, Bobo Ju, Siao Liu, Yuzheng Wang, Dingkang Yang, Peng Sun, and Liang Song. 2024b. AMP-Net: Appearance-Motion Prototype Network Assisted Automatic Video Anomaly Detection System. IEEE Transactions on Industrial Informatics 20, 2 (2024), 2843–2855.
- Liu et al. (2022c) Yang Liu, Jing Liu, Mengyang Zhao, Dingkang Yang, Xiaoguang Zhu, and Liang Song. 2022c. Learning Appearance-Motion Normality for Video Anomaly Detection. In 2022 IEEE International Conference on Multimedia and Expo (ICME). IEEE, 1–6.
- Liu et al. (2023c) Yang Liu, Zhaoyang Xia, Mengyang Zhao, Donglai Wei, Yuzheng Wang, Liu Siao, Bobo Ju, Gaoyun Fang, Jing Liu, and Liang Song. 2023c. Learning Causality-inspired Representation Consistency for Video Anomaly Detection. In Proceedings of the 31st ACM International Conference on Multimedia. 203–212.
- Liu et al. (2024c) Yang Liu, Dingkang Yang, Yan Wang, Jing Liu, Jun Liu, Azzedine Boukerche, Peng Sun, and Liang Song. 2024c. Generalized video anomaly event detection: Systematic taxonomy and comparison of deep models. Comput. Surveys (2024).
- Liu et al. (2021b) Zhian Liu, Yongwei Nie, Chengjiang Long, Qing Zhang, and Guiqing Li. 2021b. A hybrid video anomaly detection framework via memory-augmented flow reconstruction and flow-guided frame prediction. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 13588–13597.
- Liu et al. (2023b) Zuhao Liu, Xiao-Ming Wu, Dian Zheng, Kun-Yu Lin, and Wei-Shi Zheng. 2023b. Generating anomalies for video anomaly detection with prompt-based feature mapping. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 24500–24510.
- Lu et al. (2013) Cewu Lu, Jianping Shi, and Jiaya Jia. 2013. Abnormal event detection at 150 fps in matlab. In Proceedings of the IEEE international conference on computer vision. 2720–2727.
- Luo et al. (2017) Weixin Luo, Wen Liu, and Shenghua Gao. 2017. Remembering history with convolutional lstm for anomaly detection. In 2017 IEEE International Conference on Multimedia and Expo (ICME). IEEE, 439–444.
- Luo et al. (2021) Weixin Luo, Wen Liu, Dongze Lian, and Shenghua Gao. 2021. Future frame prediction network for video anomaly detection. IEEE Transactions on Pattern Analysis and Machine Intelligence (2021).
- Luo et al. (2019) Weixin Luo, Wen Liu, Dongze Lian, Jinhui Tang, Lixin Duan, Xi Peng, and Shenghua Gao. 2019. Video anomaly detection with sparse coding inspired deep neural networks. IEEE transactions on pattern analysis and machine intelligence 43, 3 (2019), 1070–1084.
- Mahadevan et al. (2010) Vijay Mahadevan, Weixin Li, Viral Bhalodia, and Nuno Vasconcelos. 2010. Anomaly detection in crowded scenes. In 2010 IEEE computer society conference on computer vision and pattern recognition. IEEE, 1975–1981.
- Mašková et al. (2024) Michaela Mašková, Matěj Zorek, Tomáš Pevnỳ, and Václav Šmídl. 2024. Deep anomaly detection on set data: Survey and comparison. Pattern Recognition (2024), 110381.
- Medel and Savakis (2016) Jefferson Ryan Medel and Andreas Savakis. 2016. Anomaly detection in video using predictive convolutional long short-term memory networks. arXiv preprint arXiv:1612.00390 (2016).
- Mishra et al. (2024) Pratik K Mishra, Alex Mihailidis, and Shehroz S Khan. 2024. Skeletal Video Anomaly Detection Using Deep Learning: Survey, Challenges, and Future Directions. IEEE Transactions on Emerging Topics in Computational Intelligence (2024).
- Nawarathna et al. (2014) Ruwan Nawarathna, JungHwan Oh, Jayantha Muthukudage, Wallapak Tavanapong, Johnny Wong, Piet C De Groen, and Shou Jiang Tang. 2014. Abnormal image detection in endoscopy videos using a filter bank and local binary patterns. Neurocomputing 144 (2014), 70–91.
- Nawaratne et al. (2019) Rashmika Nawaratne, Damminda Alahakoon, Daswin De Silva, and Xinghuo Yu. 2019. Spatiotemporal anomaly detection using deep learning for real-time video surveillance. IEEE Transactions on Industrial Informatics 16, 1 (2019), 393–402.
- Nayak et al. (2021) Rashmiranjan Nayak, Umesh Chandra Pati, and Santos Kumar Das. 2021. A comprehensive review on deep learning-based methods for video anomaly detection. Image and Vision Computing 106 (2021), 104078.
- Nguyen et al. (2020) Khac-Tuan Nguyen, Dat-Thanh Dinh, Minh N Do, and Minh-Triet Tran. 2020. Anomaly detection in traffic surveillance videos with gan-based future frame prediction. In Proceedings of the 2020 International Conference on Multimedia Retrieval. 457–463.
- Nguyen and Meunier (2019) Trong-Nguyen Nguyen and Jean Meunier. 2019. Anomaly detection in video sequence with appearance-motion correspondence. In Proceedings of the IEEE/CVF international conference on computer vision. 1273–1283.
- Ning et al. (2024) Zhiyuan Ning, Zile Wang, Yang Liu, Jing Liu, and Liang Song. 2024. Memory-enhanced appearance-motion consistency framework for video anomaly detection. Computer Communications 216 (2024), 159–167.
- Oprea et al. (2020) Sergiu Oprea, Pablo Martinez-Gonzalez, Alberto Garcia-Garcia, John Alejandro Castro-Vargas, Sergio Orts-Escolano, Jose Garcia-Rodriguez, and Antonis Argyros. 2020. A review on deep learning techniques for video prediction. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 6 (2020), 2806–2826.
- Pang et al. (2021) Guansong Pang, Chunhua Shen, Longbing Cao, and Anton Van Den Hengel. 2021. Deep learning for anomaly detection: A review. ACM Computing Surveys (CSUR) 54, 2 (2021), 1–38.
- Pang et al. (2020) Guansong Pang, Cheng Yan, Chunhua Shen, Anton van den Hengel, and Xiao Bai. 2020. Self-trained deep ordinal regression for end-to-end video anomaly detection. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 12173–12182.
- Park et al. (2020) Hyunjong Park, Jongyoun Noh, and Bumsub Ham. 2020. Learning memory-guided normality for anomaly detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 14372–14381.
- Qiu et al. (2024) Shaoming Qiu, Jingfeng Ye, Jiancheng Zhao, Lei He, Liangyu Liu, E Bicong, and Xinchen Huang. 2024. Video anomaly detection guided by clustering learning. Pattern Recognition (2024), 110550.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International conference on machine learning. PMLR, 8748–8763.
- Raja et al. (2022) Rohit Raja, Prakash Chandra Sharma, Md Rashid Mahmood, and Dinesh Kumar Saini. 2022. Analysis of anomaly detection in surveillance video: recent trends and future vision. Multimedia Tools and Applications (2022), 1–17.
- Ramachandra and Jones (2020) Bharathkumar Ramachandra and Michael Jones. 2020. Street Scene: A new dataset and evaluation protocol for video anomaly detection. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 2569–2578.
- Ramachandra et al. (2020) Bharathkumar Ramachandra, Michael J Jones, and Ranga Raju Vatsavai. 2020. A survey of single-scene video anomaly detection. IEEE transactions on pattern analysis and machine intelligence 44, 5 (2020), 2293–2312.
- Ravanbakhsh et al. (2018) Mahdyar Ravanbakhsh, Moin Nabi, Hossein Mousavi, Enver Sangineto, and Nicu Sebe. 2018. Plug-and-play cnn for crowd motion analysis: An application in abnormal event detection. In 2018 IEEE Winter Conference on Applications of Computer Vision (WACV). IEEE, 1689–1698.
- Ravanbakhsh et al. (2017) Mahdyar Ravanbakhsh, Moin Nabi, Enver Sangineto, Lucio Marcenaro, Carlo Regazzoni, and Nicu Sebe. 2017. Abnormal event detection in videos using generative adversarial nets. In 2017 IEEE international conference on image processing (ICIP). IEEE, 1577–1581.
- Ravanbakhsh et al. (2019) Mahdyar Ravanbakhsh, Enver Sangineto, Moin Nabi, and Nicu Sebe. 2019. Training adversarial discriminators for cross-channel abnormal event detection in crowds. In 2019 IEEE Winter Conference on Applications of Computer Vision (WACV). IEEE, 1896–1904.
- Reddy et al. (2021) Mr D Murahari Reddy, Mr Sk Masthan Basha, Mr M Chinnaiahgari Hari, and Mr N Penchalaiah. 2021. Dall-e: Creating images from text. UGC Care Group I Journal 8, 14 (2021), 71–75.
- Redmon et al. (2016) Joseph Redmon, Santosh Divvala, Ross Girshick, and Ali Farhadi. 2016. You only look once: Unified, real-time object detection. In Proceedings of the IEEE conference on computer vision and pattern recognition. 779–788.
- Rezaee et al. (2021) Khosro Rezaee, Sara Mohammad Rezakhani, Mohammad R Khosravi, and Mohammad Kazem Moghimi. 2021. A survey on deep learning-based real-time crowd anomaly detection for secure distributed video surveillance. Personal and Ubiquitous Computing (2021), 1–17.
- Roady et al. (2020) Ryne Roady, Tyler L Hayes, Hitesh Vaidya, and Christopher Kanan. 2020. Stream-51: Streaming classification and novelty detection from videos. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition Workshops. 228–229.
- Roshtkhari and Levine (2013) Mehrsan Javan Roshtkhari and Martin D Levine. 2013. An on-line, real-time learning method for detecting anomalies in videos using spatio-temporal compositions. Computer vision and image understanding 117, 10 (2013), 1436–1452.
- Sabokrou et al. (2015) Mohammad Sabokrou, Mahmood Fathy, Mojtaba Hoseini, and Reinhard Klette. 2015. Real-time anomaly detection and localization in crowded scenes. In Proceedings of the IEEE conference on computer vision and pattern recognition workshops. 56–62.
- Sabokrou et al. (2017) Mohammad Sabokrou, Mohsen Fayyaz, Mahmood Fathy, and Reinhard Klette. 2017. Deep-cascade: Cascading 3d deep neural networks for fast anomaly detection and localization in crowded scenes. IEEE Transactions on Image Processing 26, 4 (2017), 1992–2004.
- Saligrama and Chen (2012) Venkatesh Saligrama and Zhu Chen. 2012. Video anomaly detection based on local statistical aggregates. In 2012 IEEE Conference on computer vision and pattern recognition. IEEE, 2112–2119.
- Saligrama et al. (2010) Venkatesh Saligrama, Janusz Konrad, and Pierre-Marc Jodoin. 2010. Video anomaly identification. IEEE Signal Processing Magazine 27, 5 (2010), 18–33.
- Samaila et al. (2024) Yau Alhaji Samaila, Patrick Sebastian, Narinderjit Singh Sawaran Singh, Aliyu Nuhu Shuaibu, Syed Saad Azhar Ali, Temitope Ibrahim Amosa, Ghulam E Mustafa Abro, and Isiaka Shuaibu. 2024. Video Anomaly Detection: A Systematic Review of Issues and Prospects. Neurocomputing (2024), 127726.
- Samariya and Thakkar (2023) Durgesh Samariya and Amit Thakkar. 2023. A comprehensive survey of anomaly detection algorithms. Annals of Data Science 10, 3 (2023), 829–850.
- Santhosh et al. (2020) Kelathodi Kumaran Santhosh, Debi Prosad Dogra, and Partha Pratim Roy. 2020. Anomaly detection in road traffic using visual surveillance: A survey. ACM Computing Surveys (CSUR) 53, 6 (2020), 1–26.
- Shi et al. (2021) Ting Shi, Wu Yang, and Junfei Qiao. 2021. Research on Nonlinear Systems Modeling Methods Based on Neural Networks. In Journal of Physics: Conference Series, Vol. 2095. IOP Publishing, 012037.
- Singh and Pankajakshan (2018) Prakhar Singh and Vinod Pankajakshan. 2018. A Deep Learning Based Technique for Anomaly Detection in Surveillance Videos. In 2018 Twenty Fourth National Conference on Communications (NCC). 1–6.
- Smeureanu et al. (2017) Sorina Smeureanu, Radu Tudor Ionescu, Marius Popescu, and Bogdan Alexe. 2017. Deep appearance features for abnormal behavior detection in video. In International Conference on Image Analysis and Processing. Springer, 779–789.
- Song et al. (2022) Liang Song, Xing Hu, Guanhua Zhang, Petros Spachos, Konstantinos N Plataniotis, and Hequan Wu. 2022. Networking systems of AI: on the convergence of computing and communications. IEEE Internet of Things Journal 9, 20 (2022), 20352–20381.
- Sultani et al. (2018) Waqas Sultani, Chen Chen, and Mubarak Shah. 2018. Real-world anomaly detection in surveillance videos. In Proceedings of the IEEE conference on computer vision and pattern recognition. 6479–6488.
- Sun et al. (2019) Jiayu Sun, Jie Shao, and Chengkun He. 2019. Abnormal event detection for video surveillance using deep one-class learning. Multimedia Tools and Applications 78, 3 (2019), 3633–3647.
- Sun et al. (2022) Zehua Sun, Qiuhong Ke, Hossein Rahmani, Mohammed Bennamoun, Gang Wang, and Jun Liu. 2022. Human action recognition from various data modalities: A review. IEEE transactions on pattern analysis and machine intelligence 45, 3 (2022), 3200–3225.
- Sun et al. (2024) Zhe Sun, Panpan Wang, Wang Zheng, and Meng Zhang. 2024. Dual GroupGAN: An unsupervised Four-Competitor (2V2) approach for video anomaly detection. Pattern Recognition (2024), 110500.
- Tian et al. (2024) Beiwen Tian, Huan-ang Gao, Leiyao Cui, Yupeng Zheng, Lan Luo, Baofeng Wang, Rong Zhi, Guyue Zhou, and Hao Zhao. 2024. Latency-aware Road Anomaly Segmentation in Videos: A Photorealistic Dataset and New Metrics. arXiv preprint arXiv:2401.04942 (2024).
- Tian et al. (2021) Yu Tian, Guansong Pang, Yuanhong Chen, Rajvinder Singh, Johan W Verjans, and Gustavo Carneiro. 2021. Weakly-supervised video anomaly detection with robust temporal feature magnitude learning. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 4975–4986.
- Tran et al. (2015) Du Tran, Lubomir Bourdev, Rob Fergus, Lorenzo Torresani, and Manohar Paluri. 2015. Learning spatiotemporal features with 3d convolutional networks. In Proceedings of the IEEE international conference on computer vision. 4489–4497.
- Tran and Hogg (2017) Hanh TM Tran and David Hogg. 2017. Anomaly detection using a convolutional winner-take-all autoencoder. In Proceedings of the British Machine Vision Conference 2017. British Machine Vision Association.
- Turchini et al. (2017) Francesco Turchini, Lorenzo Seidenari, and Alberto Del Bimbo. 2017. Convex polytope ensembles for spatio-temporal anomaly detection. In International Conference on Image Analysis and Processing. Springer, 174–184.
- Vu et al. (2017) Hung Vu, Dinh Phung, Tu Dinh Nguyen, Anthony Trevors, and Svetha Venkatesh. 2017. Energy-based models for video anomaly detection. arXiv preprint arXiv:1708.05211 (2017).
- Walia et al. (2023) Guneet Kaur Walia, Mohit Kumar, and Sukhpal Singh Gill. 2023. AI-empowered fog/edge resource management for IoT applications: A comprehensive review, research challenges and future perspectives. IEEE Communications Surveys & Tutorials (2023).
- Wang et al. (2018) Tian Wang, Meina Qiao, Zhiwei Lin, Ce Li, Hichem Snoussi, Zhe Liu, and Chang Choi. 2018. Generative neural networks for anomaly detection in crowded scenes. IEEE Transactions on Information Forensics and Security 14, 5 (2018), 1390–1399.
- Wang et al. (2021) Xuanzhao Wang, Zhengping Che, Bo Jiang, Ning Xiao, Ke Yang, Jian Tang, Jieping Ye, Jingyu Wang, and Qi Qi. 2021. Robust unsupervised video anomaly detection by multipath frame prediction. IEEE transactions on neural networks and learning systems (2021).
- Wang et al. (2022) Yunbo Wang, Haixu Wu, Jianjin Zhang, Zhifeng Gao, Jianmin Wang, S Yu Philip, and Mingsheng Long. 2022. Predrnn: A recurrent neural network for spatiotemporal predictive learning. IEEE Transactions on Pattern Analysis and Machine Intelligence 45, 2 (2022), 2208–2225.
- Wu et al. (2024a) Peng Wu, Jing Liu, Xiangteng He, Yuxin Peng, Peng Wang, and Yanning Zhang. 2024a. Toward Video Anomaly Retrieval From Video Anomaly Detection: New Benchmarks and Model. IEEE Transactions on Image Processing 33 (2024), 2213–2225.
- Wu et al. (2019) Peng Wu, Jing Liu, and Fang Shen. 2019. A deep one-class neural network for anomalous event detection in complex scenes. IEEE transactions on neural networks and learning systems 31, 7 (2019), 2609–2622.
- Wu et al. (2020) Peng Wu, Jing Liu, Yujia Shi, Yujia Sun, Fangtao Shao, Zhaoyang Wu, and Zhiwei Yang. 2020. Not only look, but also listen: Learning multimodal violence detection under weak supervision. In European conference on computer vision. Springer, 322–339.
- Wu et al. (2022) Peng Wu, Xiaotao Liu, and Jing Liu. 2022. Weakly supervised audio-visual violence detection. IEEE Transactions on Multimedia (2022).
- Wu et al. (2024b) Peng Wu, Xuerong Zhou, Guansong Pang, Yujia Sun, Jing Liu, Peng Wang, and Yanning Zhang. 2024b. Open-Vocabulary Video Anomaly Detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
- Wu et al. (2024c) Peng Wu, Xuerong Zhou, Guansong Pang, Lingru Zhou, Qingsen Yan, Peng Wang, and Yanning Zhang. 2024c. Vadclip: Adapting vision-language models for weakly supervised video anomaly detection. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 6074–6082.
- Wu et al. (2010) Shandong Wu, Brian E Moore, and Mubarak Shah. 2010. Chaotic invariants of lagrangian particle trajectories for anomaly detection in crowded scenes. In 2010 IEEE computer society conference on computer vision and pattern recognition. IEEE, 2054–2060.
- Xu et al. (2017) Dan Xu, Yan Yan, Elisa Ricci, and Nicu Sebe. 2017. Detecting anomalous events in videos by learning deep representations of appearance and motion. Computer Vision and Image Understanding 156 (2017), 117–127.
- Xu et al. (2019) Hang Xu, Lewei Yao, Wei Zhang, Xiaodan Liang, and Zhenguo Li. 2019. Auto-fpn: Automatic network architecture adaptation for object detection beyond classification. In Proceedings of the IEEE/CVF international conference on computer vision. 6649–6658.
- Xu et al. (2023) Peng Xu, Xiatian Zhu, and David A Clifton. 2023. Multimodal learning with transformers: A survey. IEEE Transactions on Pattern Analysis and Machine Intelligence (2023).
- Yan et al. (2018) Shiyang Yan, Jeremy S Smith, Wenjin Lu, and Bailing Zhang. 2018. Abnormal event detection from videos using a two-stream recurrent variational autoencoder. IEEE Transactions on Cognitive and Developmental Systems 12, 1 (2018), 30–42.
- Yang et al. (2021) Jingkang Yang, Kaiyang Zhou, Yixuan Li, and Ziwei Liu. 2021. Generalized out-of-distribution detection: A survey. arXiv preprint arXiv:2110.11334 (2021).
- Yang et al. (2024) Zhiwei Yang, Jing Liu, and Peng Wu. 2024. Text Prompt with Normality Guidance for Weakly Supervised Video Anomaly Detection. arXiv preprint arXiv:2404.08531 (2024).
- Ye et al. (2024) Hongyu Ye, Ke Xu, Xinghao Jiang, and Tanfeng Sun. 2024. Learning Spatio-Temporal Relations with Multi-Scale Integrated Perception for Video Anomaly Detection. In ICASSP 2024-2024 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 4020–4024.
- Ye et al. (2019) Muchao Ye, Xiaojiang Peng, Weihao Gan, Wei Wu, and Yu Qiao. 2019. Anopcn: Video anomaly detection via deep predictive coding network. In Proceedings of the 27th ACM International Conference on Multimedia. 1805–1813.
- Yin et al. (2024) Yafeng Yin, Lei Xie, Zhiwei Jiang, Fu Xiao, Jiannong Cao, and Sanglu Lu. 2024. A Systematic Review of Human Activity Recognition Based On Mobile Devices: Overview, Progress and Trends. IEEE Communications Surveys & Tutorials (2024).
- Yu et al. (2023) Guang Yu, Siqi Wang, Zhiping Cai, Xinwang Liu, En Zhu, and Jianping Yin. 2023. Video Anomaly Detection via Visual Cloze Tests. IEEE Transactions on Information Forensics and Security (2023).
- Yu et al. (2020) Guang Yu, Siqi Wang, Zhiping Cai, En Zhu, Chuanfu Xu, Jianping Yin, and Marius Kloft. 2020. Cloze test helps: Effective video anomaly detection via learning to complete video events. In Proceedings of the 28th ACM International Conference on Multimedia. 583–591.
- Yu et al. (2021) Jongmin Yu, Younkwan Lee, Kin Choong Yow, Moongu Jeon, and Witold Pedrycz. 2021. Abnormal event detection and localization via adversarial event prediction. IEEE Transactions on Neural Networks and Learning Systems (2021).
- Zaheer et al. (2020) Muhammad Zaigham Zaheer, Jin-ha Lee, Marcella Astrid, and Seung-Ik Lee. 2020. Old is gold: Redefining the adversarially learned one-class classifier training paradigm. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 14183–14193.
- Zaheer et al. (2022) M Zaigham Zaheer, Arif Mahmood, M Haris Khan, Mattia Segu, Fisher Yu, and Seung-Ik Lee. 2022. Generative Cooperative Learning for Unsupervised Video Anomaly Detection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 14744–14754.
- Zhang et al. (2022) Qianqian Zhang, Guorui Feng, and Hanzhou Wu. 2022. Surveillance video anomaly detection via non-local U-Net frame prediction. Multimedia Tools and Applications (2022), 1–16.
- Zhang et al. (2015) Zhenzhen Zhang, Jianjun Hou, Qinglong Ma, and Zhaohong Li. 2015. Efficient video frame insertion and deletion detection based on inconsistency of correlations between local binary pattern coded frames. Security and Communication networks 8, 2 (2015), 311–320.
- Zhao et al. (2022) Mengyang Zhao, Yang Liu, Jing Liu, and Xinhua Zeng. 2022. Exploiting Spatial-temporal Correlations for Video Anomaly Detection. In 2022 26th International Conference on Pattern Recognition (ICPR). IEEE, 1727–1733.
- Zhao et al. (2017) Yiru Zhao, Bing Deng, Chen Shen, Yao Liu, Hongtao Lu, and Xian-Sheng Hua. 2017. Spatio-temporal autoencoder for video anomaly detection. In Proceedings of the 25th ACM international conference on Multimedia. 1933–1941.
- Zhou et al. (2019) Joey Tianyi Zhou, Jiawei Du, Hongyuan Zhu, Xi Peng, Yong Liu, and Rick Siow Mong Goh. 2019. Anomalynet: An anomaly detection network for video surveillance. IEEE Transactions on Information Forensics and Security 14, 10 (2019), 2537–2550.
- Zhou et al. (2016) Shifu Zhou, Wei Shen, Dan Zeng, Mei Fang, Yuanwang Wei, and Zhijiang Zhang. 2016. Spatial–temporal convolutional neural networks for anomaly detection and localization in crowded scenes. Signal Processing: Image Communication 47 (2016), 358–368.
- Zhu et al. (2022) Yuansheng Zhu, Wentao Bao, and Qi Yu. 2022. Towards open set video anomaly detection. In European Conference on Computer Vision. Springer, 395–412.